### v3

# The 2D free particle in the phase space quantum mechanics

Abstract
The Wigner eigenfunctions of a free quantum particle propagating on a plane are derived. Two possibilities are analysed. First, the particle with fixed components of the Cartesian momentum is considered. Then the particle of given energy and angular momentum is discussed. In that second case, a special choice of coordinates on the symplectic space $(\mathbb{R}^{4},\,\omega)$ suitable for the representation of eigenstates of the discussed particle is presented. Further, the Moyal $\star_{(\text{M})}$-product on the phase space is derived with the use of the Fedosov algorithm adapted to these coordinates on a flat phase space. Next, the eigenvalue equations for the Wigner eigenfunction are solved and the physically acceptable solutions are identified. Finally, a relationship between the Wigner eigenfunction of the particle with the fixed components of the Cartesian momentum and the cross-Wigner functions of the particle with the given energy and angular momentum is found.
1 Introduction
The phase space quantum mechanics  [1, 2, 3, 4, 5, 6, 7, 8, 9] is an interesting alternative to the Hilbert space version of quantum theory. Its power manifests when applied to systems living in nontrivial configuration spaces but it can be fruitful also in more elementary problems.
For physical systems, which classically are modelled on the symplectic manifold $(\mathbb{R}^{2n},\omega),$ the quantum phase space is the same as the classical one. And quantum observables are represented by smooth real-valued functions on that manifold. On the contrary, states are not identified with points on the phase space but with quasiprobability distributions known as Wigner functions. Moreover, the multiplication of quantum observables is no longer the standard pointwise multiplication $\cdot$’ of functions. Instead, it is given by an associative, bilinear operation $\star_{(\text{M})}$’, known as the Moyal product.
The eigenvalue equations for observables are usually partial differential equations for Wigner eigenfunctions. Such equations may be difficult to solve. However, applying e.g. differential geometry would make this process easier.
In one of our earlier works [10] we presented a method for solving the $\star_{(\text{M})}$-eigenvalue problem for a 1D Hamiltonian. In that publication we built the star eigenvalue equation in some convenient Darboux coordinates by constructing the $\star_{(\text{M})}$-product with the use of the Fedosov algorithm [11, 12].
In the current article, we extend that approach to solve the star-eigenvalue problem for a free quantum particle of mass $M$ moving in the space $\mathbb{R}^{2}$. The particle is characterised either by components of the Cartesian momentum or by energy and the angular momentum.
The first option is easily solvable. The second case is much more demanding. In order to deal with it we choose some special Darboux coordinates $(T,\chi,H,L)$. Then employing the adapted Fedosov algorithm, we explicitly construct the $\star_{(\text{M})}$-product in the new chart. We formulate the $\star_{(\text{M})}$-eigenvalue problem for both the Hamiltonian $H$ and the angular momentum $L$ and solve two $\star_{(\text{M})}$-eigenvalue problems explicitly, obtaining a collection of Wigner eigenfunctions $W_{Em}.$ Finally, we examine which of the resulting functions represent physically acceptable Wigner eigenfunctions of the 2D free particle of given energy $E$ and the magnetic quantum number $m$.
The last part of our work refers to the vital question of how to decompose states in combinations of some quasiprobability distributions. Namely, knowing the Wigner eigenfunction $W_{p_{x0}p_{y0}}$ for the 2D free molecule with the fixed components of the Cartesian momentum $p_{x0},p_{y0}$ and some cross-Wigner functions $W_{Emm^{\prime}}$ for the 2D free particle with the defined energy $E$ and angular momenta $m\hbar,m^{\prime}\hbar$ we find an expansion of $W_{p_{x0}p_{y0}}$ in the generalised linear combination of functions from the set $\{W_{Emm^{\prime}}\}_{E>0,m,\,m^{\prime}\in\mathbb{Z}}.$ This result illustrates the structure of set of states as a linear space.
2 Wigner eigenfunctions of Cartesian momentum and energy–angular momentum
Let us consider a free nonrelativistic classical particle of mass $M$ moving in $\mathbb{R}^{2}$ space. Its phase space is the symplectic manifold $(\mathbb{R}^{4},\,\omega).$ One can cover it with the Cartesian chart $(x,\,y,\,p_{x},\,p_{y})$ in which the symplectic form equals
$\omega=\text{d}x\wedge\text{d}p_{x}+\text{d}y\wedge\text{d}p_{y}.$ (2.1)
The motion of this classical particle is determined by three independent constants of motion: the components of the Cartesian momentum $p_{x},p_{y}$ and the angular momentum $L=xp_{y}-yp_{x}.$
Eigenstates of a free quantum particle living on the plane $\mathbb{R}^{2}$ are characterised either by values of components of the momentum $p_{x},p_{y}$ or the energy being value of the Hamilton function $H$ and the angular momentum $L.$ These two kinds of eigenstates are not equivalent i.e. a quantum particle with defined components of the momentum has an undetermined angular momentum and vice versa.
Let us discuss the aforementioned eigenstates of the 2D free particle of mass $M$ in frames of phase space quantum mechanics. The quantum phase space of this particle is—like in the classical physics—the symplectic manifold $(\mathbb{R}^{4},\,\omega).$
We need to introduce the Moyal product on the symplectic manifold. This product is a counterpart of composition of linear operators in a Hilbert space. The Moyal product $\star_{(\text{M})}$ of functions $f(\mathbf{r},\,\mathbf{p})$ and $g(\mathbf{r},\,\mathbf{p})$ is defined (compare [18, 17]) in the Cartesian chart $(\mathbf{r},\,\mathbf{p})\equiv(x,\,y,\,p_{x},\,p_{y})$ as
$(f\star_{(\text{M})}g)(\mathbf{r},\,\mathbf{p}):=\frac{1}{(\pi\hbar)^{4}}\int_{\mathbb{R}^{8}}\text{d}\mathbf{r}^{\prime}\text{d}\mathbf{p}^{\prime}\text{d}\mathbf{r}^{\prime\prime}\text{d}\mathbf{p}^{\prime\prime}$
$\times f(\mathbf{r}^{\prime},\mathbf{p}^{\prime})g(\mathbf{r}^{\prime\prime},\mathbf{p}^{\prime\prime})\exp\Bigg(\frac{2i}{\hbar}[(\mathbf{r}-\mathbf{r}^{\prime})(\mathbf{p}-\mathbf{p}^{\prime\prime})-(\mathbf{r}-\mathbf{r}^{\prime\prime})(\mathbf{p}-\mathbf{p}^{\prime})]\Bigg).$ (2.2)
In several cases however, one can apply the differential formula
$(f\star_{(\text{M})}g)(\mathbf{r},\,\mathbf{p})$ $:=f(\mathbf{r},\,\mathbf{p})\;\text{exp}\left(\frac{i\hbar}{2}\left[\frac{\overleftarrow{\partial}}{\partial x}\frac{\overrightarrow{\partial}}{\partial p_{x}}-\frac{\overleftarrow{\partial}}{\partial p_{x}}\frac{\overrightarrow{\partial}}{\partial x}+\frac{\overleftarrow{\partial}}{\partial y}\frac{\overrightarrow{\partial}}{\partial p_{y}}-\frac{\overleftarrow{\partial}}{\partial p_{y}}\frac{\overrightarrow{\partial}}{\partial y}\right]\right)g(\mathbf{r},\,\mathbf{p}).$ (2.3)
The Moyal bracket $\{\cdot,\,\cdot\}_{(\text{M})}$ of functions $f(\mathbf{r},\,\mathbf{p})$ and $g(\mathbf{r},\,\mathbf{p})$ being a counterpart of the commutator of respective operators is defined as
$\{f,\,g\}_{(\text{M})}(\mathbf{r},\,\mathbf{p}):=\frac{1}{i\hbar}\big(f\star_{(\text{M})}g-g\star_{(\text{M})}f\big)(\mathbf{r},\,\mathbf{p}).$ (2.4)
As we already mentioned, eigenstates of quantum 2D free particle can be determined by two complete sets of observables $\{R_{1},\,R_{2}\}$ commuting in the sense of the Moyal bracket:
  1. (i)
    the components of the momentum: $R_{1}:=p_{x}$ and $R_{2}:=p_{y}$,
  2. (ii)
    the energy and the angular momentum: $R_{1}:=H=\frac{p^{2}_{x}+p^{2}_{y}}{2M}$ and $R_{2}:=L=xp_{y}-yp_{x}$.
These two sets of observables are neither equivalent nor complementary, because
$\{p_{x},\,L\}_{(\text{M})}\neq 0\;\;\;,\;\;\;\{p_{y},\,L\}_{(\text{M})}\neq 0.$
The Wigner eigenfunction $W_{r_{1}r_{2}}(\mathbf{r},\,\mathbf{p})$ of the 2D free particle is a real function or a tempered distribution on $\mathbb{R}^{4}$ that satisfies the system of eigenvalue equations
$(R_{1}\star_{(\text{M})}W_{r_{1}r_{2}})(\mathbf{r},\,\mathbf{p})=r_{1}W_{r_{1}r_{2}}(\mathbf{r},\,\mathbf{p}),$ (2.5a)
$(R_{2}\star_{(\text{M})}W_{r_{1}r_{2}})(\mathbf{r},\,\mathbf{p})=r_{2}W_{r_{1}r_{2}}(\mathbf{r},\,\mathbf{p})$ (2.5b)
together with the commutativity conditions
$\{R_{1},\,W_{r_{1}r_{2}}\}_{(\text{M})}(\mathbf{r},\,\mathbf{p})$ = 0, (2.5c)
$\{R_{2},\,W_{r_{1}r_{2}}\}_{(\text{M})}(\mathbf{r},\,\mathbf{p})$ = 0. (2.5d)
By $r_{1}$ and $r_{2}$ we denote the eigenvalues of functions $R_{1}$ and $R_{2}$ respectively. Moreover, the Wigner eigenfunctions $W_{r_{1}r_{2}}(\mathbf{r},\,\mathbf{p})$ have to obey the condition of positivity of the marginal distributions
$\int_{\mathbb{R}^{2}}\text{d}\mathbf{r}W_{r_{1}r_{2}}(\mathbf{r},\,\mathbf{p})$ $\geq 0$ , $\int_{\mathbb{R}^{2}}\text{d}\mathbf{p}W_{r_{1}r_{2}}(\mathbf{r},\,\mathbf{p})$ $\geq 0$ (2.6)
and some normalization condition which explicit form will be derived in the next section
$(W_{r_{1}r_{2}}\star_{(\text{M})}W_{r^{\prime}_{1}r^{\prime}_{2}})(\mathbf{r},\,\mathbf{p})\propto\frac{1}{(2\pi\hbar)^{2}}f_{r_{1}r^{\prime}_{1}}f_{r_{2}r^{\prime}_{2}}W_{r_{1}r_{2}}(\mathbf{r},\,\mathbf{p}),$ (2.7)
where for i=1, 2i=1,\,2
[…]
We do not expect that any of Wigner eigenfunctions $W_{r_{1}r_{2}}(\mathbf{r},\,\mathbf{p})$ satisfies the condition $\int_{\mathbb{R}^{4}}\text{d}\mathbf{r}\text{d}\mathbf{p}W_{r_{1}r_{2}}(\mathbf{r},\,\mathbf{p})<\infty$ as the particle of concern is free.
In the case when components of the momentum of particle are determined, the solution of Eqs. (2.5a)–(2.5d) is well known
$W_{p_{x0}p_{y0}}(x,\,y,\,p_{x},\,p_{y})=\frac{1}{(2\pi\hbar)^{2}}\delta(p_{x}-p_{x0})\delta(p_{y}-p_{y0}).$ (2.9)
Symbols $p_{x0},\,p_{y0}\in\mathbb{R}$ denote eigenvalues of the Cartesian components of the momentum. One can check that the Wigner eigenfunction (2.9) also satisfies conditions (2.6).
Let us consider the second option—the 2D particle with the defined energy and angular momentum. One can see that $R_{1}=H$ and $R_{2}=L$ are quadratic functions of position and momentum. Hence, the Moyal brackets in Eqs. (2.5c)–(2.5d) are proportional to the respective Poisson brackets. This means that if we find coordinates $T$ and $\chi$ such that the transformation $(x,\,y,\,p_{x},p_{y})\to(T,\,\chi,\,H,\,L)$ is symplectic, then the LHS of Eqs. (2.5c)–(2.5d) reveals its simplest form: $\{\cdot,\,H\}_{(\text{M})}=\frac{\partial}{\partial T}(\cdot)$ and $\{\cdot,\,L\}_{(\text{M})}=\frac{\partial}{\partial\chi}(\cdot)$. In order to derive the explicit expressions of $T$ and $\chi$, one considers conservation of the symplectic form $\omega$ under the aforementioned transformation. One obtains that
$T$ $=\frac{M(xp_{x}+yp_{y})}{p^{2}_{x}+p^{2}_{y}},$ $\chi$ $=\text{atan2}(p_{y},\,p_{x}),$
$H$ $=\frac{p^{2}_{x}+p^{2}_{y}}{2M},$ $L$ $=xp_{y}-yp_{x}.$ (2.10)
The transformation (2.10) is well-defined only if $p_{x}^{2}+p_{y}^{2}>0$ i.e. it does not work for the classical particles at rest. Fortunately, the subset $p_{x}^{2}+p^{2}_{y}=0$ is a set of measure zero. Hence, it can be ignored from a quantum point of view. One can check that the Jacobian of the transformation equals
$\frac{\partial(T,\,\chi,\,H,\,L)}{\partial(x,\,y,\,p_{x},\,p_{y})}=1.$ (2.11)
The classical interpretation of the variable $T$ is the following. It is the time of arrival for the particle of mass $M$ moving with the velocity $\frac{\mathbf{p}}{M}$ from the origin (0, 0)(0,\,0) to the point being the projection of the point $\mathbf{r}$} on the direction of the momentum $\mathbf{p}.$
Assuming $p_{x}^{2}+p_{y}^{2}>0$, the inverse transformation is given
$x$ $=\frac{2HT\cos\chi+L\sin\chi}{\sqrt{2MH}},$ $y$ $=\frac{2HT\sin\chi-L\cos\chi}{\sqrt{2MH}},$
$p_{x}$ $=\sqrt{2MH}\cos\chi,$ $p_{y}$ $=\sqrt{2MH}\sin\chi.$ (2.12)
Now, we would like to know how the Moyal product (2.3) looks in the new coordinates (2.10). The most efficient way to achieve this is to apply the Fedosov algorithm  [11, 12] which requires equipping our phase space $(\mathbb{R}^{4},\,\omega)$ with a symplectic connection $\gamma$. The triple $(\mathbb{R}^{4},\,\omega,\gamma)$ is known as the Fedosov manifold [13]. We assume that in the Darboux chart $(x,\,y,\,p_{x},\,p_{y})$ all the symplectic connection $\gamma$ coefficients disappear
$\gamma^{i}_{\,\,\,jk}$ $=0,$ $i,\,j,\,k$ $=1,\,2,\,3,\,4.$ (2.13)
Under the change of coordinates
$(Q^{a}):=(x,\,y,\,p_{x},\,p_{y})\to(\widetilde{Q}^{a}):=(T,\,\chi,\,H,\,L),$
due to the condition (2.13), the symplectic connection coefficients $\gamma_{ijk}$ transform according to the rule [14]
$\widetilde{\gamma}_{ijk}$ $=\omega_{rd}\frac{\partial Q^{r}}{\partial\widetilde{Q}^{i}}\frac{\partial^{2}Q^{d}}{\partial\widetilde{Q}^{j}\partial\widetilde{Q}^{k}},$ $i,\,j,\,k$ $=1,\,2,\,3,\,4.$ (2.14)
In expression (2.14) the Einstein summation convention is used. The coefficients of symplectic connection $\widetilde{\gamma}_{ijk}$ are symmetric in all indices. Thus, the only non-zero representatives of the symplectic connection are
$\widetilde{\gamma}_{122}$ $=-2H,$ $\widetilde{\gamma}_{133}$ $=-\frac{1}{2H},$ $\widetilde{\gamma}_{222}$ $=-2\,{L},$ $\widetilde{\gamma}_{233}$ $=\frac{L}{2H^{2}},$ $\widetilde{\gamma}_{234}$ $=-\frac{1}{2H}.$ (2.15)
We apply the Fedosov algorithm (see Appendix A) to construct the Fedosov star product $\star$’ for the symplectic connection (2.15). In this case the star product $\star$’ is simply the Moyal product $\star_{(\text{M})}$’ in the coordinates (2.10). Then, using the already constructed $\star$-product in Eqs. (2.5a)–(2.5d) for $R_{1}=H$ and $R_{2}=L$ we receive that the Wigner eigenfunction $W_{Em}(T,\,\chi,\,H,\,L)$ of energy $E>0$ and of angular momentum $m\hbar$, where $m\in\mathbb{R}$, fulfills the system of equations
$HW_{Em}-\frac{\hbar^{2}}{4}\bigg(H\frac{\partial^{2}W_{Em}}{\partial L^{2}}+\frac{1}{4H}\frac{\partial^{2}W_{Em}}{\partial T^{2}}\bigg)=EW_{Em},$ (2.16a)
$LW_{Em}-\frac{\hbar^{2}}{4}\bigg(2\frac{\partial W_{Em}}{\partial L}+L\frac{\partial^{2}W_{Em}}{\partial L^{2}}+2H\frac{\partial^{2}W_{Em}}{\partial H\partial L}+\frac{1}{4H}\frac{\partial^{2}W_{Em}}{\partial T\partial\chi}$
$-\frac{L}{4H^{2}}\frac{\partial^{2}W_{Em}}{\partial T^{2}}\bigg)=m\hbar W_{Em}$ (2.16b)
with the additional conditions
$\frac{\partial W_{Em}}{\partial T}$ = 0, (2.16c)
$\frac{\partial W_{Em}}{\partial\chi}$ = 0. (2.16d)
The general solution of Eqs. (2.16c)–(2.16d) is a function depending exclusively on the variables $H$ and $L$. Therefore the Eqs. (2.16a)–(2.16b) turn into
$HW_{Em}-\frac{\hbar^{2}}{4}H\frac{\partial^{2}W_{Em}}{\partial L^{2}}=EW_{Em},$ (2.17a)
$LW_{Em}-\frac{\hbar^{2}}{2}\left(\frac{\partial W_{Em}}{\partial L}+H\frac{\partial^{2}W_{Em}}{\partial L\partial H}+\frac{L}{2}\frac{\partial^{2}W_{Em}}{\partial L^{2}}\right)=m\hbar W_{Em}.$ (2.17b)
In the case when $H>E$ one can show that the general solution of Eq. (2.17a) is of the form
$W_{Em}(T,\,\chi,\,H,\,L)=A_{1}(H)\text{exp}\left(\frac{2L}{\hbar}\sqrt{\frac{H-E}{H}}\right)+A_{2}(H)\text{exp}\left(-\frac{2L}{\hbar}\sqrt{\frac{H-E}{H}}\right).$ (2.18)
The functions $A_{1}(H)$ and $A_{2}(H)$ are arbitrary but real. The absolute value of solution (2.18) increases rapidly for $L\rightarrow\pm\infty$ so it cannot be a Wigner eigenfunction. Therefore in the region $H>E$ the only one physically acceptable solution is
$W_{Em}(T,\,\chi,\,H,\,L)=0.$ (2.19)
In the case when $0<H<E$ the general real solution of Eq. (2.17a) is of the form
$W_{Em}(T,\,\chi,\,H,\,L)=B_{1}(H)\cos\bigg(\frac{2L}{\hbar}\sqrt{\frac{E-H}{H}}\bigg)+B_{2}(H)\sin\bigg(\frac{2L}{\hbar}\sqrt{\frac{E-H}{H}}\bigg),$ (2.20)
where $B_{1}(H)$ and $B_{2}(H)$ are arbitrary real functions of $H$. Substituting the solution (2.20) into Eq. (2.17b), we obtain a differential equation for the functions $B_{1}(H)$ and $B_{2}(H)$. Solving it, we receive that the complete solution of (2.17a)–(2.17b) for all $H>0$ is of the form
$W_{Em}(T,\,\chi,\,H,\,L)=\frac{Y(E-H)N_{Em}}{\sqrt{H(E-H)}}\cos\left(\frac{2L}{\hbar}\sqrt{\frac{E-H}{H}}-2m\arccos\sqrt{\frac{H}{E}}+D_{Em}\right),$ (2.21)
where $N_{Em},\,D_{Em}$ are some real constants which may depend on $E$ and $m$. The symbol $Y(x)$ denotes the Heaviside step function
$Y(x):=\begin{cases}1,&x>0,\\ 0,&x<0.\end{cases}$ (2.22)
It turns out that function (2.21) satisfies Eq. (2.17b) in a distributional sense only if $D_{Em}=a\pi$, where $a\in\mathbb{Z}$. Let us put $D_{Em}=0$. Thus, for all $E>0$ and $m\in\mathbb{R}$, we obtain
$W_{Em}(T,\,\chi,\,H,\,L)=\frac{Y\left(E-H\right)N_{Em}}{\sqrt{H(E-H)}}\cos\left(\frac{2L}{\hbar}\sqrt{\frac{E-H}{H}}-2m\arccos\sqrt{\frac{H}{E}}\right).$ (2.23)
Notice that the function $W_{Em}$ written in the form (2.23) depends exclusively on the variables $H$ and $L$ representing energy and classical angular momentum.
Lets prove that the solution (2.23) is the Wigner eigenfunction only if $m\in\mathbb{Z}$. In order to do that we use the coordinates $(r,\,\phi,\,p,\,\chi)$, where the pair $(r,\,\phi)$ denotes the polar coordinates of the position vector $\mathbf{r}$ and $(p,\,\chi)$ are the polar coordinates of the momentum vector $\mathbf{p}$. These new coordinates are not symplectic. The elementary volume $\text{d}T\text{d}\chi\text{d}H\text{d}L$ transforms into $r\text{d}r\text{d}\phi p\text{d}p\text{d}\chi.$
The energy is of the form $E=\frac{p^{2}_{0}}{2M}$, where $p_{0}$ denotes the length of some momentum vector. Then the function (2.23) turns into
$W_{Em}(r,\,\phi,\,p,\,\chi)=\frac{Y(p_{0}-p)2MN_{Em}}{p\sqrt{p_{0}^{2}-p^{2}}}\cos\left(\frac{2r\sqrt{p_{0}^{2}-p^{2}}}{\hbar}\sin(\chi-\phi)-2m\arccos{\frac{p}{p_{0}}}\right).$ (2.24)
Denote the integrals for any arbitrary fixed $m\in\mathbb{R}$ as
$P_{Em}(r,\,\phi):=\int^{\infty}_{0}p\text{d}p\int^{2\pi}_{0}\text{d}\chi W_{Em}(r,\,\phi,\,p,\,\chi).$ (2.25)
If $W_{Em}$ is a Wigner eigenfunction, the integral (2.25) represents the spatial density of probability. Substituting (2.24) to (2.25) and integrating over $\chi$, we obtain that
$P_{Em}(r,\,\phi)$ $=4\pi MN_{Em}\int^{p_{0}}_{0}\text{d}p\frac{1}{\sqrt{p^{2}_{0}-p^{2}}}\cos\left(2m\arccos\frac{p}{p_{0}}\right)J_{0}\left(\frac{2r\sqrt{p^{2}_{0}-p^{2}}}{\hbar}\right).$ (2.26)
By $J_{n}(x),\;n\in\mathbb{Z}$ we mean the Bessel function of the first kind. One can see that function $P_{Em}$ does not depend on $\phi$. We introduce $\theta$ fulfilling the conditions
$p=p_{0}\cos\theta,\;\;\;\theta\in\left(0,\frac{\pi}{2}\right)$ (2.27)
and receive
$P_{Em}(r,\,\phi)$ $=4\pi MN_{Em}\int^{\frac{\pi}{2}}_{0}\text{d}\theta\cos(2m\theta)J_{0}\left(\frac{2p_{0}r\sin\theta}{\hbar}\right).$ (2.28)
We are going to discuss two cases for the magnetic number value $m$: non-integer and integer.
Firstly, we examine the case when $m\in\mathbb{Z}$. Then, the integral (2.28) simplifies to
$P_{Em}(r,\,\phi)=2\pi^{2}MN_{Em}J^{2}_{m}\left(\frac{p_{0}r}{\hbar}\right).$ (2.29)
Moreover, we can determine the normalization constant $N_{Em}$ by postulating the normalization condition (2.7) for $m,\,\widetilde{m}\in\mathbb{Z}$
$(W_{Em}\star_{\text{(M)}}W_{\widetilde{E}\widetilde{m}})(T,\,\chi,\,H,\,L)=\frac{1}{(2\pi\hbar)^{2}M}\delta(E-\widetilde{E})\delta_{m\widetilde{m}}W_{Em}(T,\,\chi,\,H,\,L).$ (2.30)
Considering the condition (2.30) and the result of Appendix B i.e.
$\big(W_{Em}\star_{(\text{M})}W_{\widetilde{E}m}\big)(T,\,\chi,\,H,\,L)=\pi N_{Em}\delta(E-\widetilde{E})W_{Em}(T,\,\chi,\,H,\,L),$ (2.31)
we obtain
$N_{Em}=\frac{1}{4\pi^{3}M\hbar^{2}}.$ (2.32)
Hence, the solution written in the form (2.24) with the normalization constant (2.32) coincides with the result presented in [15].
Let us assume $m\in\mathbb{R}\setminus\mathbb{Z}$. Unfortunately, in that case, we do not know the closed form of the integral (2.28). We choose the same $N_{Em}$ as (2.32) and do numerical calculations. The figure 1 illustrates that the integral (2.28) for a non-integer quantum angular momentum number $m$ becomes negative for some radii $r$. It contradicts the interpretation of $P_{Em}(r,\,\phi)$ as a spatial density of probability.
Figure 1: The numerical integral $P_{Em}(r,\,\phi)$, given by the formula (2.28), as a function of radius $r$ for energy $E=100\,\mathrm{eV}$, mass $M=9\cdot 10^{-31}\,\mathrm{kg}$ and three non-integer quantum magnetic numbers $m$. The normalization constant $N_{Em}$ is chosen as (2.32).
At Fig. 2 we show two examples of the Wigner eigenfunctions $W_{Em}.$ One can see that they take both positive and negative values as expected. For the fixed angular momentum $L\neq 0$ and the variable $H\to 0^{+}$ the limit of function $W_{Em}$ does not exist. However, if $H\to E^{-}$, then for arbitrary fixed $L$ the limit of $W_{Em}$ equals $\infty$.
Figure 2: The Wigner eigenfunction $W_{Em}(T,\,\chi,\,H,\,L)$, given by the formula (2.23) for energy $E=100\,\mathrm{eV}$, mass $M=9\cdot 10^{-31}\,\mathrm{kg}$ and magnetic quantum number (a) $m=0$, (b) $m=5$. The normalization constant $N_{Em}$ equals (2.32).
3 The Wigner eigenfunction of Cartesian momentum expanded in the cross-Wigner functions of energy-angular momentum
In this section, we find an expansion of the Wigner eigenfunction of Cartesian momentum in the cross-Wigner functions of energy-angular momentum. From physical point of view, it is problem of superposition in phase space quantum mechanics. In the Hilbert space quantum mechanics, such expansion is known as the Jacobi-Anger expansion.
In the previous chapter, we have derived the Wigner eigenfunction $W_{Em}(\mathbf{r},\,\mathbf{p})$, given by Eq. (2.23). In fact, it represents the state $|\Psi_{Em}\rangle$ with a given energy $E>0$ and angular momentum $m\hbar$, where $m\in\mathbb{Z}$. Analogously, the Wigner eigenfunction $W_{p_{x0}p_{y0}}(\mathbf{r},\,\mathbf{p})$ (see Eq. (2.9)) represents the state $|\Psi_{p_{x0}p_{y0}}\rangle$ with a given Cartesian momentum $(p_{x0},\,p_{y0})\in\mathbb{R}^{2}$. We will denote the cross Wigner function of the above states $|\Psi_{Em}\rangle$ and $|\Psi_{Em^{\prime}}\rangle$ as $W_{Emm^{\prime}}(\mathbf{r},\,\mathbf{p})$. It is defined for each $E>0,\,m,m^{\prime}\in\mathbb{Z}$ as a solution of the system of star-eigenvalue equations
$(H\star_{(\text{M})}W_{Emm^{\prime}})(\mathbf{r},\,\mathbf{p})=EW_{Emm^{\prime}}(\mathbf{r},\,\mathbf{p}),$ (3.1a)
$(L\star_{(\text{M})}W_{Emm^{\prime}})(\mathbf{r},\,\mathbf{p})=m\hbar W_{Emm^{\prime}}(\mathbf{r},\,\mathbf{p}),$ (3.1b)
$(W_{Emm^{\prime}}\star_{(\text{M})}L)(\mathbf{r},\,\mathbf{p})=m^{\prime}\hbar W_{Emm^{\prime}}(\mathbf{r},\,\mathbf{p})$ (3.1c)
fulfilling the commutation rule
$\{H,\,W_{Emm^{\prime}}\}_{(\text{M})}(\mathbf{r},\,\mathbf{p})$ $=0$ (3.1d)
and satisfying the generalised normalization condition
$\big(W_{Emm^{\prime}}\star_{(\text{M})}W_{\widetilde{E}\widetilde{m}\widetilde{m}^{\prime}}\big)(T,\,\chi,\,H,\,L)$ $=\frac{1}{(2\pi\hbar)^{2}M}\delta(E-\widetilde{E})\delta_{m^{\prime}\widetilde{m}}W_{Em\widetilde{m}^{\prime}}(T,\,\chi,\,H,\,L).$ (3.1e)
Thus we do not demand the functions $W_{Emm^{\prime}}(\mathbf{r},\,\mathbf{p})$ to be real. From Eqs. (3.1b)–(3.1c) and the property $\overline{A\star_{\text{(M)}}B}=\overline{B}\star_{\text{(M)}}\overline{A}$ one can derive that
$W_{Emm^{\prime}}(\mathbf{r},\,\mathbf{p})=\overline{W}_{Em^{\prime}m}(\mathbf{r},\,\mathbf{p}).$ (3.2)
In the coordinates $(T,\,\chi,\,H,\,L)$ defined by the relations (2.10), the formulas (3.1a) and (3.1d) look like Eqs. (2.16a) and (2.16c). Therefore we postulate that
$W_{Emm^{\prime}}(T,\,\chi,\,H,\,L)=$
$Y(E-H)\bigg(B_{1}(H,\,\chi)\cos\bigg(\frac{2L}{\hbar}\sqrt{\frac{E-H}{H}}\bigg)+B_{2}(H,\,\chi)\sin\bigg(\frac{2L}{\hbar}\sqrt{\frac{E-H}{H}}\bigg)\bigg).$ (3.3)
In order to find functions $B_{1}(H,\,\chi)$ and $B_{2}(H,\,\chi)$, we substitute the solution (3.3) into Eqs. (3.1b) and (3.1c). Then, using linear independence of $\sin\bigg(\frac{2L}{\hbar}\sqrt{\frac{E-H}{H}}\bigg)$ and $\cos\bigg(\frac{2L}{\hbar}\sqrt{\frac{E-H}{H}}\bigg)$, one can deduce that
$(m-m^{\prime})B_{1}+i\frac{\partial B_{1}}{\partial\chi}=0,$ (3.4a)
$(m-m^{\prime})B_{2}+i\frac{\partial B_{2}}{\partial\chi}=0,$ (3.4b)
$\sqrt{(E-H)H}(m+m^{\prime})B_{1}+(E-2H)B_{2}+2(E-H)H\frac{\partial B_{2}}{\partial H}=0,$ (3.4c)
$-\sqrt{(E-H)H}(m+m^{\prime})B_{2}+(E-2H)B_{1}+2(E-H)H\frac{\partial B_{1}}{\partial H}=0.$ (3.4d)
Thus finally
$W_{Emm^{\prime}}(T,\,\chi,\,H,\,L)=\frac{Y\left(E-H\right)N_{Emm^{\prime}}}{\sqrt{H(E-H)}}e^{i(m-m^{\prime})\chi}$
$\times\cos\left(\frac{2L}{\hbar}\sqrt{\frac{E-H}{H}}-(m+m^{\prime})\arccos\sqrt{\frac{H}{E}}\right).$ (3.5)
Here $N_{Emm^{\prime}}$ is a normalization constant that satisfies the relation (3.1e) and also the condition
$N_{Emm^{\prime}}=\overline{N}_{Em^{\prime}m},$ (3.6)
due to Eq. (3.2).
In Appendix B we have proved that
$\big(W_{Emm^{\prime}}\star_{(\text{M})}W_{\widetilde{E}m^{\prime}\widetilde{m}^{\prime}}\big)(T,\,\chi,\,H,\,L)=\frac{\pi N_{Emm^{\prime}}N_{\widetilde{E}m^{\prime}\widetilde{m}^{\prime}}}{N_{Em\widetilde{m}^{\prime}}}\delta(E-\widetilde{E})W_{Em\widetilde{m}^{\prime}}(T,\,\chi,\,H,\,L).$ (3.7)
Comparing (3.1e) and (3.7), we deduce that
$N_{Emm^{\prime}}$ $=\frac{1}{4\pi^{3}M\hbar^{2}}z^{m-m^{\prime}},$ (3.8)
where $z\in\mathbb{C}\setminus\{0\}$ is such that Eq. (3.6) holds i.e. $z$ has to satisfy the requirement
$z^{m-m^{\prime}}=\overline{z^{m^{\prime}-m}}.$ (3.9)
Using the property $|z^{n}|=|z|^{n}$ for $n\in\mathbb{Z}$, Eq. (3.9) induces that $|z|=1$. Therefore, the normalization constant can be determined up to a phase factor $\alpha\in\mathbb{R}$
$N_{Emm^{\prime}}=\frac{1}{4\pi^{3}M\hbar^{2}}\exp(i(m-m^{\prime})\alpha).$ (3.10)
We want to find such coefficients $C_{\widetilde{m}\widetilde{m}^{\prime}}$ that
$W_{\widetilde{p}_{x0}\widetilde{p}_{y0}}(T,\,\chi,\,H,\,L)=\sum^{\infty}_{\widetilde{m}=-\infty}\sum^{\infty}_{\widetilde{m}^{\prime}=-\infty}C_{\widetilde{m}\widetilde{m}^{\prime}}W_{\widetilde{E}\widetilde{m}\widetilde{m}^{\prime}}(T,\,\chi,\,H,\,L),$ (3.11)
where the Wigner eigenfunction $W_{\widetilde{p}_{x0}\widetilde{p}_{y0}}$ is defined as (2.9). In the coordinates (2.10), this function looks like
$W_{\widetilde{p}_{x0}\widetilde{p}_{y0}}(T,\,\chi,\,H,\,L)=N_{\widetilde{p}_{x0}\widetilde{p}_{y0}}\delta(H-\widetilde{E})\delta(\chi-\widetilde{\chi}_{0}).$ (3.12)
The coefficients in the formula (3.12) are equal to
$N_{\widetilde{p}_{x0}\widetilde{p}_{y0}}$ $:=\frac{1}{(2\pi\hbar)^{2}M},$ $\widetilde{E}$ $:=\frac{\widetilde{p}^{2}_{x0}+\widetilde{p}^{2}_{y0}}{2M},$ $\widetilde{\chi}_{0}$ $:=\text{atan2}(\widetilde{p}_{y0},\,\widetilde{p}_{x0}).$ (3.13)
Expression (3.11) represents the state $|\Psi_{\widetilde{p}_{x0}\widetilde{p}_{y0}}\rangle$ as a superposition of states $\{|\Psi_{\widetilde{E}m}\rangle\}_{m\in\mathbb{Z}}$ with $\widetilde{E}=\frac{\widetilde{p}^{2}_{x0}+\widetilde{p}^{2}_{y0}}{2M}$. We multiply both sides of Eq. (3.11) by $W_{Emm^{\prime}}$ in the sense of the Moyal product
$(W_{Emm^{\prime}}\star_{(\text{M})}W_{\widetilde{p}_{x0}\widetilde{p}_{y0}})(T,\,\chi,\,H,\,L)=\underbrace{\sum^{\infty}_{\widetilde{m}=-\infty}\sum^{\infty}_{\widetilde{m}^{\prime}=-\infty}C_{\widetilde{m}\widetilde{m}^{\prime}}(W_{Emm^{\prime}}\star_{(\text{M})}W_{\widetilde{E}\widetilde{m}\widetilde{m}^{\prime}})(T,\,\chi,\,H,\,L)}_{=:Q}.$ (3.14)
In Appendix C, we have shown that the LHS of Eq. (3.14) equals
$\big(W_{Emm^{\prime}}\star_{(\text{M})}W_{\widetilde{p}_{x0}\widetilde{p}_{y0}}\big)(T,\,\chi,\,H,\,L)=N_{Emm^{\prime}}N_{\widetilde{p}_{x0}\widetilde{p}_{y0}}\frac{Y\big(\cos(\chi-\widetilde{\chi}_{0})\big)}{\cos^{2}(\chi-\widetilde{\chi}_{0})}e^{2im\chi}e^{-i(m+m^{\prime})\widetilde{\chi}_{0}}$
$\times\exp\bigg(-\frac{2i}{\hbar}L\tan(\chi-\widetilde{\chi}_{0})\bigg)\delta\bigg(E-\widetilde{E}\bigg)\delta\bigg(E-\frac{H}{\cos^{2}(\chi-\widetilde{\chi}_{0})}\bigg).$ (3.15)
Let us focus on the RHS of Eq. (3.14), denoted as $Q$. We substitute the normalization condition (3.1e), the normalization constant (3.13) and the explicit form of $W_{Emm^{\prime}}(\mathbf{r},\,\mathbf{p})$ given by Eq. (3.5). We postulate that
$C_{\widetilde{m}\widetilde{m}^{\prime}}:=\frac{1}{2\pi}\frac{N_{Em\widetilde{m}}}{N_{Em\widetilde{m}^{\prime}}}e^{i(-\widetilde{m}+\widetilde{m}^{\prime})\widetilde{\chi}_{0}}=\frac{1}{2\pi}e^{i(-\widetilde{m}+\widetilde{m}^{\prime})\alpha}e^{i(-\widetilde{m}+\widetilde{m}^{\prime})\widetilde{\chi}_{0}}.$ (3.16)
Then
$Q=N_{Emm^{\prime}}N_{\widetilde{p}_{x0}\widetilde{p}_{y0}}\delta(E-\widetilde{E})\sum^{\infty}_{\widetilde{m}^{\prime}=-\infty}\frac{1}{2\pi}e^{i(-m^{\prime}+\widetilde{m}^{\prime})\widetilde{\chi}_{0}}$
$\times\frac{Y(E-H)}{\sqrt{H(E-H)}}e^{i(m-\widetilde{m}^{\prime})\chi}\cos\bigg(\frac{2L}{\hbar}\sqrt{\frac{E-H}{H}}-(m+\widetilde{m}^{\prime})\arccos\sqrt{\frac{H}{E}}\bigg).$ (3.17)
Using the formula $\cos(x)=\frac{e^{ix}+e^{-ix}}{2}$, one can write $Q$ as
$Q=N_{Emm^{\prime}}N_{\widetilde{p}_{x0}\widetilde{p}_{y0}}\delta(E-\widetilde{E})\frac{1}{4\pi}\frac{Y(E-H)}{\sqrt{H(E-H)}}e^{i(m\chi-m^{\prime}\widetilde{\chi}_{0})}$
$\times\bigg\{\exp\bigg(i\frac{2L}{\hbar}\sqrt{\frac{E-H}{H}}\bigg)\exp\bigg(-im\arccos\sqrt{\frac{H}{E}}\bigg)\sum^{\infty}_{\widetilde{m}^{\prime}=-\infty}\exp\bigg(i\widetilde{m}^{\prime}\bigg[-\chi+\widetilde{\chi}_{0}-\arccos\sqrt{\frac{H}{E}}\bigg]\bigg)$
$+\exp\bigg(-i\frac{2L}{\hbar}\sqrt{\frac{E-H}{H}}\bigg)\exp\bigg(im\arccos\sqrt{\frac{H}{E}}\bigg)\sum^{\infty}_{\widetilde{m}^{\prime}=-\infty}\exp\bigg(i\widetilde{m}^{\prime}\bigg[-\chi+\widetilde{\chi}_{0}+\arccos\sqrt{\frac{H}{E}}\bigg]\bigg)\bigg\}.$
Using the identity $\sum^{\infty}_{n=-\infty}e^{inx}=2\pi\sum^{\infty}_{k=-\infty}\delta(x+2k\pi)$ for $x\in\mathbb{R}$, we receive
$Q=N_{Emm^{\prime}}N_{\widetilde{p}_{x0}\widetilde{p}_{y0}}\delta(E-\widetilde{E})\frac{1}{2}\frac{Y(E-H)}{\sqrt{H(E-H)}}e^{i(m\chi-m^{\prime}\widetilde{\chi}_{0})}$
$\times\bigg\{\exp\bigg(i\frac{2L}{\hbar}\sqrt{\frac{E-H}{H}}\bigg)\exp\bigg(-im\arccos\sqrt{\frac{H}{E}}\bigg)\sum^{\infty}_{k=-\infty}\delta\bigg(-\chi+\widetilde{\chi}_{0}-\arccos\sqrt{\frac{H}{E}}+2k\pi\bigg)$
$+\exp\bigg(-i\frac{2L}{\hbar}\sqrt{\frac{E-H}{H}}\bigg)\exp\bigg(im\arccos\sqrt{\frac{H}{E}}\bigg)\sum^{\infty}_{k=-\infty}\delta\bigg(-\chi+\widetilde{\chi}_{0}+\arccos\sqrt{\frac{H}{E}}+2k\pi\bigg)\bigg\}.$
The Dirac deltas lead to the following set of constraints
$\chi+\widetilde{\chi}_{0}\mp\arccos\sqrt{\frac{H}{E}}+2k\pi=0.$ (3.18)
The solution of Eq. (3.18) exists only if:
  1. (i)
    $\chi-\widetilde{\chi}_{0}\in(3\pi/2,\,2\pi)\text{mod}(2\pi)$ for the ‘-’ choice,
  2. (ii)
    $\chi-\widetilde{\chi}_{0}\in(0,\,\pi/2)\text{mod}(2\pi)$ for the ‘++’ choice.
In both cases, the solution of Eq. (3.18) is of the form
$E=\frac{H}{\cos^{2}\big(\chi-\widetilde{\chi}_{0}\big)}.$ (3.19)
Then, using the Dirac delta coordinates transformation rule with respect to $E$, we obtain that
$\sum^{\infty}_{k=-\infty}\delta\bigg($ $-\chi+\widetilde{\chi}_{0}\mp\arccos\sqrt{\frac{H}{E}}+2k\pi\bigg)$
$=2E\sqrt{\frac{E-H}{H}}\delta\bigg(E-\frac{H}{\cos^{2}\big(\chi-\widetilde{\chi}_{0}\big)}\bigg)Y\big(\cos\big(\chi-\widetilde{\chi}_{0}\big)\big)Y(\mp\sin(\chi-\widetilde{\chi}_{0})).$ (3.20)
Therefore
$Q=N_{Emm^{\prime}}N_{\widetilde{p}_{x0}\widetilde{p}_{y0}}\delta(E-\widetilde{E})\frac{E}{H}Y(E-H)e^{i(m\chi-m^{\prime}\widetilde{\chi}_{0})}Y\big(\cos\big(\chi-\widetilde{\chi}_{0}\big)\big)\delta\bigg(E-\frac{H}{\cos^{2}\big(\chi-\widetilde{\chi}_{0}\big)}\bigg)$
$\times\bigg\{\exp\bigg(i\frac{2L}{\hbar}\sqrt{\frac{E-H}{H}}\bigg)\exp\bigg(-im\arccos\sqrt{\frac{H}{E}}\bigg)Y(-\sin(\chi-\widetilde{\chi}_{0}))$
$+\exp\bigg(-i\frac{2L}{\hbar}\sqrt{\frac{E-H}{H}}\bigg)\exp\bigg(im\arccos\sqrt{\frac{H}{E}}\bigg)Y(+\sin(\chi-\widetilde{\chi}_{0}))\bigg\}.$ (3.21)
Finally, we can use the identity $f(x)\delta(x-a)=f(a)\delta(x-a)$ in order to retrieve Eq. (3.15). It proves that
$W_{\widetilde{p}_{x0}\widetilde{p}_{y0}}(T,\,\chi,\,H,\,L)=\sum^{\infty}_{\widetilde{m}=-\infty}\sum^{\infty}_{\widetilde{m}^{\prime}=-\infty}\frac{1}{2\pi}e^{i(-\widetilde{m}+\widetilde{m}^{\prime})\alpha}e^{i(-\widetilde{m}+\widetilde{m}^{\prime})\widetilde{\chi}_{0}}W_{\widetilde{E}\widetilde{m}\widetilde{m}^{\prime}}(T,\,\chi,\,H,\,L).$ (3.22)
In the Hilbert space quantum mechanics, the Jacobi-Anger expansion [19] holds for the wavefunctions $\psi_{Em}(r,\,\phi)=\frac{1}{\sqrt{2\pi}\hbar}J_{m}(k_{0}r)e^{im\phi}$ (compare with [15, 20]) and $\psi_{p_{x0}p_{y0}}(r,\,\phi)=\frac{1}{2\pi\hbar}e^{irk_{0}\cos(\chi_{0}-\phi)}$ with $E=\frac{(\hbar k_{0})^{2}}{2M}$
$\psi_{p_{x0}p_{y0}}(r,\,\phi)=\sum^{\infty}_{m=-\infty}\frac{1}{\sqrt{2\pi}}i^{m}e^{-im\chi_{0}}\psi_{Em}(r,\,\phi).$ (3.23)
Thus, for $\alpha=-\pi/2$, the expansion (3.22) is compatible with Eq. (3.23).
Statement
The work was partially financed under the program ‘FU2N - Fund for the Improvement of Skills of Young Scientists’ supporting the research excellence of employees and doctoral students of the Lodz University of Technology.
Acknowledgment
This article has been completed while the first author was the Doctoral Candidate in the Interdisciplinary Doctoral School at the Lodz University of Technology, Poland.
Appendix A An adaptation of the construction of the Fedosov star product to symplectic flat manifolds
This Appendix contains a sketch of the algorithm [11, 12] of constructing a star product on an arbitrary Fedosov manifold. Some parts of the original Fedosov proposal are modified, since we deal exclusively with flat symplectic manifolds. As before, the Einstein summation convention is applied.
Let $(\mathcal{M},\,\omega,\gamma)$ be a $2n$-dimensional flat Fedosov manifold [13] i.e. a $2n$-D manifold $\mathcal{M}$ equipped with a symplectic structure $\omega$ and a flat connection $\gamma$ which is torsion free and which preserves the symplectic form $\omega$.
The Fedosov algorithm works only in the Darboux coordinates $(x^{1},\ldots,x^{2n})$, but every symplectic manifold can be covered with a Darboux atlas. In these coordinates the matrix $\omega_{ij}$ of symplectic form and its inverse $\omega^{ij}$ are equal respectively
$(\omega_{ij})$ $=\begin{bmatrix}\mathbb{O}&\mathbb{I}\\ -\mathbb{I}&\mathbb{O}\end{bmatrix}$, $(\omega^{ij})$ $=\begin{bmatrix}\mathbb{O}&-\mathbb{I}\\ \mathbb{I}&\mathbb{O}\end{bmatrix}$, (A.1)
where the symbol $\mathbb{O}$ denotes the zero square matrix of dimension $n\times n$ and the symbol $\mathbb{I}$ represents the $n\times n$ identity matrix.
The symplectic connection $\gamma$ is determined by the coefficients $\gamma^{i}_{\,\,\,jk}$ or equivalently by $\gamma_{ijk}:=\omega_{ir}\gamma^{r}_{\,\,\,jk}$. In every Darboux chart $(x^{1},\ldots,x^{2n})$ the coefficients $\gamma_{ijk}$ are symmetric in all indices
$\gamma_{ijk}$ $=\gamma_{ikj}=\gamma_{jik}=\gamma_{jki}=\gamma_{kij}=\gamma_{kji},$ $i,\,j,\,k$ $=1,\,\dots,\,2n.$ (A.2)
[…]
$W_{x}:=\left\{a=\sum^{\infty}_{k,\,l=0}\hbar^{k}a_{k,i_{1}\dots i_{l}}y^{i_{1}}\dots y^{i_{l}}:(y^{1},\,\dots,\,y^{2n})\in T_{x}\mathcal{M},\,a_{k,i_{1}\dots i_{l}}\in(T^{*}_{x}\mathcal{M})^{\otimes_{s}l}\right\}$ (A.3)
and for which their $\circ$-product equals
$a\circ b:=\sum^{\infty}_{k=0}\left(\frac{i\hbar}{2}\right)^{k}\frac{1}{k!}\omega^{i_{1}j_{1}}\dots\omega^{i_{k}j_{k}}\frac{\partial^{k}a}{\partial y^{i_{1}}\dots\partial y^{i_{k}}}\frac{\partial^{k}b}{\partial y^{j_{1}}\dots\partial y^{j_{k}}}.$ (A.4)
[…]
$W:=\bigcup_{x\in\mathcal{M}}W_{x}$ (A.5)
with the fibrewise multiplication (A.4).
[…]
$W\otimes\Lambda^{q}:=\Bigg\{\sum^{\infty}_{k,\,l=0}\hbar^{k}a_{k,\,i_{1}\dots i_{l},j_{1}\dots j_{q}}(x^{1},\ldots,x^{2n})y^{i_{1}}\dots y^{i_{l}}\text{d}x^{j_{1}}\wedge\ldots\wedge\text{d}x^{j_{q}}:\forall_{x\in\mathcal{M}}$
$a_{k,\,i_{l}\dots i_{l},j_{1}\dots j_{q}}(x^{1},\ldots,x^{2n})\in(T^{*}_{x}\mathcal{M})^{\otimes_{s}l}\otimes(T^{*}_{x}\mathcal{M})^{\otimes_{a}q},\,(y^{1},\,\dots,\,y^{2n})\in T_{x}\mathcal{M}\Bigg\}.$ (A.6)
From now we will omit variables $(x^{1},\ldots,x^{2n})$ on the symplectic manifold $\mathcal{M}.$
The multiplication $\bullet$’ is defined by means of the exterior product of differentials $\text{d}x^{i}$ and the Weyl product (A.4)
$a\bullet b:=\Bigg($ $\sum^{\infty}_{k,\,l_{1}=0}\hbar^{k}a_{k,\,i_{1}\dots i_{l_{1}},j_{1}\dots j_{q_{1}}}y^{i_{1}}\dots y^{i_{l_{1}}}\Bigg)\circ\Bigg(\sum^{\infty}_{k,\,l_{2}=0}\hbar^{k}b_{k,\,I_{1}\dots I_{l_{2}},J_{1}\dots J_{q_{2}}}y^{I_{1}}\dots y^{I_{l_{2}}}\Bigg)$
$\times\text{d}x^{j_{1}}\wedge\ldots\wedge\text{d}x^{j_{q_{1}}}\wedge\text{d}x^{J_{1}}\wedge\ldots\wedge\text{d}x^{J_{q_{2}}}.$ (A.7)
The commutator $\cdot$$\cdot$] for any $a\in W\otimes\Lambda^{q_{1}}$ and $b\in W\otimes\Lambda^{q_{2}}$ is given by the formula
$a,\,b]:=a\bullet b-(-1)^{q_{1}q_{2}}b\bullet a.$ (A.8)
The projection $P:W\otimes\Lambda^{q}\to Z\otimes\Lambda^{q}$ assigns to any $a\in W\otimes\Lambda^{q}$ its part not containing $y^{i}$’s
$P(a):=\sum^{\infty}_{k=0}\hbar^{k}a_{k,j_{1}\dots j_{q}}\text{d}x^{j_{1}}\wedge\ldots\wedge\text{d}x^{j_{q}}.$ (A.9)
By $Z$ we mean the linear space of formal series over the ring of smooth functions $C^{\infty}({\cal M})$ in $\hbar$.
The symplectic connection $\partial:W\otimes\Lambda^{q}\to W\otimes\Lambda^{q+1}$ acts as
$\partial a:=\text{d}a+\frac{i}{\hbar}[\Gamma,\,a],$ (A.10)
where $\text{d}:=\text{d}x^{i}\wedge\frac{\partial}{\partial x^{i}}$ and $\Gamma:=\frac{1}{2}\gamma_{ijk}y^{i}y^{j}\text{d}x^{k}$.
In the next step we introduce an Abelian connection $\text{D}:W\otimes\Lambda^{q}\to W\otimes\Lambda^{q+1}$ related to the symplectic connection $\gamma_{ijk}$ and given by the formula
$\text{D}a:=\partial a+\frac{i}{\hbar}[\omega_{ij}y^{i}\text{d}x^{j},\,a].$ (A.11)
The subalgebra $W_{\text{D}}\subset W$ of the Weyl algebra corresponding to the Abelian connection $D$ is defined as a collection of flat sections
$W_{\text{D}}:=\{a\in W:\text{D}a=0\}.$ (A.12)
One can show that the element $a\in W_{\text{D}}$ assigned to $P(a)\in Z$ is given uniquely as a solution of the recurrence equation
$a=P(a)+\delta^{-1}\partial a.$ (A.13)
We stress that the formula (A.13) is valid only for flat symplectic connections.
The symbol $\delta^{-1}$ denotes a linear operator such that for each monomial
$y^{i_{1}}\dots y^{i_{l}}\text{d}x^{j_{1}}\wedge\ldots\wedge\text{d}x^{j_{q}}$ we obtain
$\delta^{-1}(y^{i_{1}}\dots y^{i_{l}}\text{d}x^{j_{1}}\wedge\ldots\wedge\text{d}x^{j_{q}}):=$
$\begin{cases}\frac{1}{l+q}\sum^{q}_{r=1}(-1)^{r+1}y^{j_{r}}y^{i_{1}}\dots y^{i_{l}}\text{d}x^{j_{1}}\wedge\ldots\overset{\vee}{\text{d}x^{j_{r}}}\ldots\wedge\text{d}x^{j_{q}}&,\,l+q>0,\\ 0&,\,l+q=0.\end{cases}$ (A.14)
One can notice that the map $\sigma:W_{\text{D}}\to Z$ defined for any $a\in W_{\text{D}}$ as
$\sigma(a):=P(a)$ (A.15)
is bijective. Thus the inverse map $\sigma^{-1}$ exists and for any $f\in Z$ is determined by the recurrence equation
$\sigma^{-1}(f)=f+\delta^{-1}\partial\sigma^{-1}(f).$ (A.16)
Therefore the Fedosov star product $\star:Z\times Z\to Z$ is constructed for each $f,\,g\in Z$ as
$f\star g:=\sigma(\sigma^{-1}(f)\circ\sigma^{-1}(g))$ (A.17)
and on the manifold ${\cal M}={\mathbb{R}}^{2n}$ is equal to the Moyal product $f\star_{(\text{M})}g$ of functions $f$ and $g.$
Appendix B Calculation of the Moyal product of functions
$(W_{Emm^{\prime}}\star_{(\text{M})}W_{\widetilde{E}\widetilde{m}\widetilde{m}^{\prime}})(T,\,\chi,\,H,\,L)$
The Moyal product of cross-Wigner functions is of the form
$\big(W_{Emm^{\prime}}\star_{(\text{M})}W_{\widetilde{E}\widetilde{m}\widetilde{m}^{\prime}}\big)(T,\,\chi,\,H,\,L)=\frac{N_{Emm^{\prime}}N_{\widetilde{E}\widetilde{m}\widetilde{m}^{\prime}}}{(\pi\hbar)^{4}}\int^{\infty}_{-\infty}\text{d}T^{\prime}\int^{2\pi}_{0}\text{d}\chi^{\prime}\int^{\infty}_{0}\text{d}H^{\prime}\int^{\infty}_{-\infty}\text{d}L^{\prime}$
$\times\int^{\infty}_{-\infty}\text{d}T^{\prime\prime}\int^{2\pi}_{0}\text{d}\chi^{\prime\prime}\int^{\infty}_{0}\text{d}H^{\prime\prime}\int^{\infty}_{-\infty}\text{d}L^{\prime\prime}\frac{Y(E-H^{\prime})}{\sqrt{H^{\prime}(E-H^{\prime})}}\frac{Y(\widetilde{E}-H^{\prime\prime})}{\sqrt{H^{\prime\prime}(\widetilde{E}-H^{\prime\prime})}}e^{i(m-m^{\prime})\chi^{\prime}}e^{i(\widetilde{m}-\widetilde{m}^{\prime})\chi^{\prime\prime}}$
$\times\cos\left(\frac{2L^{\prime}}{\hbar}\sqrt{\frac{E-H^{\prime}}{H^{\prime}}}-(m+m^{\prime})\arccos\sqrt{\frac{H^{\prime}}{E}}\right)$
[…]
$\times\exp\bigg(\frac{2i}{\hbar}\bigg\{2T\bigg[\sqrt{HH^{\prime}}\cos(\chi-\chi^{\prime})-\sqrt{HH^{\prime\prime}}\cos(\chi-\chi^{\prime\prime})\bigg]$
$+2T^{\prime}\bigg[\sqrt{H^{\prime}H^{\prime\prime}}\cos(\chi^{\prime}-\chi^{\prime\prime})-\sqrt{HH^{\prime}}\cos(\chi-\chi^{\prime})\bigg]$
$+2T^{\prime\prime}\bigg[\sqrt{HH^{\prime\prime}}\cos(\chi-\chi^{\prime\prime})-\sqrt{H^{\prime}H^{\prime\prime}}\cos(\chi^{\prime}-\chi^{\prime\prime})\bigg]$
$+L\bigg[\sqrt{\frac{H^{\prime}}{H}}\sin(\chi-\chi^{\prime})-\sqrt{\frac{H^{\prime\prime}}{H}}\sin(\chi-\chi^{\prime\prime})\bigg]+L^{\prime}\bigg[\sqrt{\frac{H}{H^{\prime}}}\sin(\chi-\chi^{\prime})+\sqrt{\frac{H^{\prime\prime}}{H^{\prime}}}\sin(\chi^{\prime}-\chi^{\prime\prime})\bigg]$
$+L^{\prime\prime}\bigg[\sqrt{\frac{H^{\prime}}{H^{\prime\prime}}}\sin(\chi^{\prime}-\chi^{\prime\prime})-\sqrt{\frac{H}{H^{\prime\prime}}}\sin(\chi-\chi^{\prime\prime})\bigg]\bigg\}\bigg).$ (B.1)
First, we make the linear change of variables $\chi^{\prime}\to\chi^{\prime}-\chi$ and $\chi^{\prime\prime}\to\chi^{\prime\prime}-\chi$. Next, we integrate over $T^{\prime}$ and $T^{\prime\prime}$, applying the identities $\int^{\infty}_{-\infty}\text{d}xe^{ikx}=2\pi\delta(k)$, $\delta(ax)=\delta(x)/|a|$ and $f(x)\delta(x-b)=f(b)\delta(x-b)$, where $a\in\mathbb{R}\setminus\{0\}$ and $b\in\mathbb{R}$. Thus we obtain
$\big(W_{Emm^{\prime}}\star_{(\text{M})}W_{\widetilde{E}\widetilde{m}\widetilde{m}^{\prime}}\big)(T,\,\chi,\,H,\,L)=\frac{N_{Emm^{\prime}}N_{\widetilde{E}\widetilde{m}\widetilde{m}^{\prime}}}{4(\pi\hbar)^{2}}e^{i(m-m^{\prime}+\widetilde{m}-\widetilde{m}^{\prime})\chi}\int^{-\chi+2\pi}_{-\chi}\text{d}\chi^{\prime}\int^{\infty}_{0}\text{d}H^{\prime}$
$\times\int^{\infty}_{-\infty}\text{d}L^{\prime}\int^{-\chi+2\pi}_{-\chi}\text{d}\chi^{\prime\prime}\int^{\infty}_{0}\text{d}H^{\prime\prime}\int^{\infty}_{-\infty}\text{d}L^{\prime\prime}\frac{Y(E-H^{\prime})}{\sqrt{H^{\prime}(E-H^{\prime})}}\frac{Y(\widetilde{E}-H^{\prime\prime})}{\sqrt{H^{\prime\prime}(\widetilde{E}-H^{\prime\prime})}}e^{i(m-m^{\prime})\chi^{\prime}}e^{i(\widetilde{m}-\widetilde{m}^{\prime})\chi^{\prime\prime}}$
$\times\cos\left(\frac{2L^{\prime}}{\hbar}\sqrt{\frac{E-H^{\prime}}{H^{\prime}}}-(m+m^{\prime})\arccos\sqrt{\frac{H^{\prime}}{E}}\right)$
[…]
$\times\exp\bigg(\frac{2i}{\hbar}\bigg\{L\bigg[-\sqrt{\frac{H^{\prime}}{H}}\sin(\chi^{\prime})+\sqrt{\frac{H^{\prime\prime}}{H}}\sin(\chi^{\prime\prime})\bigg]$
$+L^{\prime}\bigg[-\sqrt{\frac{H}{H^{\prime}}}\sin(\chi^{\prime})+\sqrt{\frac{H^{\prime\prime}}{H^{\prime}}}\sin(\chi^{\prime}-\chi^{\prime\prime})\bigg]+L^{\prime\prime}\bigg[\sqrt{\frac{H^{\prime}}{H^{\prime\prime}}}\sin(\chi^{\prime}-\chi^{\prime\prime})+\sqrt{\frac{H}{H^{\prime\prime}}}\sin(\chi^{\prime\prime})\bigg]\bigg\}\bigg)$
$\times\delta\bigg(\sqrt{H^{\prime}H^{\prime\prime}}\cos(\chi^{\prime}-\chi^{\prime\prime})-\sqrt{HH^{\prime}}\cos(\chi^{\prime})\bigg)\delta\bigg(\sqrt{H^{\prime}H^{\prime\prime}}\cos(\chi^{\prime}-\chi^{\prime\prime})-\sqrt{HH^{\prime\prime}}\cos(\chi^{\prime\prime})\bigg).$ (B.2)
Let us recall the identity [16]
$\iint_{D}\text{d}x_{1}\dots\text{d}x_{n}\delta(f_{1}(x_{1},\,\dots,\,x_{n}))\dots\delta(f_{n}(x_{1},\,\dots,\,x_{n}))h(x_{1},\,\dots,\,x_{n})$
$=\sum_{(x_{10},\,\dots,\,x_{n0})\in D_{0}}\bigg|\frac{\partial(f_{1},\,\dots,\,f_{n})}{\partial(x_{1},\,\dots,\,x_{n})}\bigg|^{-1}_{(x_{10},\,\dots,\,x_{n0})}h(x_{10},\,\dots,\,x_{n0}),$ (B.3)
where $D_{0}:=\{(x_{1},\,\dots,\,x_{n})\in D:f_{1}(x_{1},\,\dots,\,x_{n})=0\land\dots\land f_{n}(x_{1},\,\dots,\,x_{n})=0\}$.
As $H,H^{\prime}$ and $H^{\prime\prime}$ are positive, all of $\cos(\chi^{\prime}-\chi^{\prime\prime}),\cos(\chi^{\prime}),\cos(\chi^{\prime\prime})$ must be different from $0.$ Therefore the Dirac delta constraints in (B.2) are satisfied by
$H^{\prime}$ $=H\frac{\cos^{2}(\chi^{\prime\prime})}{\cos^{2}(\chi^{\prime}-\chi^{\prime\prime})}$ $\land$ $H^{\prime\prime}$ $=H\frac{\cos^{2}(\chi^{\prime})}{\cos^{2}(\chi^{\prime}-\chi^{\prime\prime})}$ (B.4)
under two extra conditions
$Y\bigg(\frac{\cos(\chi^{\prime\prime})}{\cos(\chi^{\prime}-\chi^{\prime\prime})}\bigg)=Y\bigg(\frac{\cos(\chi^{\prime})}{\cos(\chi^{\prime}-\chi^{\prime\prime})}\bigg)=1.$ (B.5)
Observe that (B.5) implies the estimation
$\cos(\chi^{\prime}-\chi^{\prime\prime})\geq\sin(\chi^{\prime})\sin(\chi^{\prime\prime}).$
One can also notice that neither $\chi^{\prime}$ nor $\chi^{\prime\prime}$ equals $\pm\pi.$
We integrate over $H^{\prime}$ and $H^{\prime\prime}$, using the formula (B.3). Hence
$\big(W_{Emm^{\prime}}\star_{(\text{M})}W_{\widetilde{E}\widetilde{m}\widetilde{m}^{\prime}}\big)(T,\,\chi,\,H,\,L)=\frac{N_{Emm^{\prime}}N_{\widetilde{E}\widetilde{m}\widetilde{m}^{\prime}}}{(\pi\hbar)^{2}}e^{i(m-m^{\prime}+\widetilde{m}-\widetilde{m}^{\prime})\chi}$
$\times\int^{-\chi+2\pi}_{-\chi}\text{d}\chi^{\prime}\int^{\infty}_{-\infty}\text{d}L^{\prime}\int^{-\chi+2\pi}_{-\chi}\text{d}\chi^{\prime\prime}\int^{\infty}_{-\infty}\text{d}L^{\prime\prime}\frac{1}{\cos^{2}(\chi^{\prime}-\chi^{\prime\prime})}e^{i(m-m^{\prime})\chi^{\prime}}e^{i(\widetilde{m}-\widetilde{m}^{\prime})\chi^{\prime\prime}}$
$\frac{\times Y\left(E-H\frac{\cos^{2}(\chi^{\prime\prime})}{\cos^{2}(\chi^{\prime}-\chi^{\prime\prime})}\right)\,Y\left(\widetilde{E}-H\frac{\cos^{2}(\chi^{\prime})}{\cos^{2}(\chi^{\prime}-\chi^{\prime\prime})}\right)Y\bigg(\frac{\cos(\chi^{\prime})}{\cos(\chi^{\prime}-\chi^{\prime\prime})}\bigg)Y\bigg(\frac{\cos(\chi^{\prime\prime})}{\cos(\chi^{\prime}-\chi^{\prime\prime})}\bigg)}{{\sqrt{\frac{H\cos^{2}(\chi^{\prime\prime})}{\cos^{2}(\chi^{\prime}-\chi^{\prime\prime})}}\sqrt{\,E-\frac{H\cos^{2}(\chi^{\prime\prime})}{\cos^{2}(\chi^{\prime}-\chi^{\prime\prime})}}}{\sqrt{\frac{H\cos^{2}(\chi^{\prime})}{\cos^{2}(\chi^{\prime}-\chi^{\prime\prime})}}\sqrt{\,\widetilde{E}-\frac{H\cos^{2}(\chi^{\prime})}{\cos^{2}(\chi^{\prime}-\chi^{\prime\prime})}}}}$
$\times\cos\left(\frac{2L^{\prime}}{\hbar}\sqrt{\frac{E\cos^{2}(\chi^{\prime}-\chi^{\prime\prime})-H\cos^{2}(\chi^{\prime\prime})}{H\cos^{2}(\chi^{\prime\prime})}}-(m+m^{\prime})\,\arccos\left(\sqrt{\frac{H}{E}}\frac{\cos(\chi^{\prime\prime})}{\cos(\chi^{\prime}-\chi^{\prime\prime})}\right)\right)$
$\times\cos\left(\frac{2L^{\prime\prime}}{\hbar}\sqrt{\frac{\widetilde{E}\cos^{2}(\chi^{\prime}-\chi^{\prime\prime})-H\cos^{2}(\chi^{\prime})}{H\cos^{2}(\chi^{\prime})}}-(\widetilde{m}+\widetilde{m}^{\prime})\,\arccos\left(\sqrt{\frac{H}{\widetilde{E}}}\frac{\cos(\chi^{\prime})}{\cos(\chi^{\prime}-\chi^{\prime\prime})}\right)\right)$
$\times\exp\bigg(\frac{2i}{\hbar}\bigg\{-L\tan(\chi^{\prime}-\chi^{\prime\prime})-L^{\prime}\tan(\chi^{\prime\prime})+L^{\prime\prime}\tan(\chi^{\prime})\bigg\}\bigg).$ (B.6)
Observe that the following identity holds
$\int^{\infty}_{-\infty}\text{d}x\cos(ax-b)e^{itx}=\pi(e^{-ib}\delta(t+a)+e^{ib}\delta(t-a)).$ (B.7)
Using Eq. (B.7), we integrate over $L^{\prime}$ and obtain
$\big(W_{Emm^{\prime}}\star_{(\text{M})}W_{\widetilde{E}\widetilde{m}\widetilde{m}^{\prime}}\big)(T,\,\chi,\,H,\,L)=\frac{N_{Emm^{\prime}}N_{\widetilde{E}\widetilde{m}\widetilde{m}^{\prime}}}{2\pi\hbar}e^{i(m-m^{\prime}+\widetilde{m}-\widetilde{m}^{\prime})\chi}\frac{1}{H}\int^{-\chi+2\pi}_{-\chi}\text{d}\chi^{\prime}$
$\times\int^{-\chi+2\pi}_{-\chi}\text{d}\chi^{\prime\prime}\int^{\infty}_{-\infty}\text{d}L^{\prime\prime}\exp\bigg(\frac{2i}{\hbar}\bigg\{-L\tan(\chi^{\prime}-\chi^{\prime\prime})+L^{\prime\prime}\tan(\chi^{\prime})\bigg\}\bigg)e^{i(m-m^{\prime})\chi^{\prime}}e^{i(\widetilde{m}-\widetilde{m}^{\prime})\chi^{\prime\prime}}$
$\times Y\left(E-H\frac{\cos^{2}(\chi^{\prime\prime})}{\cos^{2}(\chi^{\prime}-\chi^{\prime\prime})}\right)\,Y\left(\widetilde{E}-H\frac{\cos^{2}(\chi^{\prime})}{\cos^{2}(\chi^{\prime}-\chi^{\prime\prime})}\right)Y\bigg(\frac{\cos(\chi^{\prime})}{\cos(\chi^{\prime}-\chi^{\prime\prime})}\bigg)Y\bigg(\frac{\cos(\chi^{\prime\prime})}{\cos(\chi^{\prime}-\chi^{\prime\prime})}\bigg)$
$\times\frac{\cos^{2}(\chi^{\prime}-\chi^{\prime\prime})}{\cos(\chi^{\prime})\cos(\chi^{\prime\prime})}\frac{1}{\sqrt{\,E\cos^{2}(\chi^{\prime}-\chi^{\prime\prime})-H\cos^{2}(\chi^{\prime\prime})}\sqrt{\,\widetilde{E}\cos^{2}(\chi^{\prime}-\chi^{\prime\prime})-H\cos^{2}(\chi^{\prime})}}$
$\times\cos\left(\frac{2L^{\prime\prime}}{\hbar}\sqrt{\frac{\widetilde{E}\cos^{2}(\chi^{\prime}-\chi^{\prime\prime})-H\cos^{2}(\chi^{\prime})}{H\cos^{2}(\chi^{\prime})}}-(\widetilde{m}+\widetilde{m}^{\prime})\,\arccos\left(\sqrt{\frac{H}{\widetilde{E}}}\frac{\cos(\chi^{\prime})}{\cos(\chi^{\prime}-\chi^{\prime\prime})}\right)\right)$
$\times\exp\bigg(-i\,\text{sgn}(\tan(\chi^{\prime\prime}))(m+m^{\prime})\,\arccos\bigg\{\sqrt{\frac{H}{E}}\frac{\cos(\chi^{\prime\prime})}{\cos(\chi^{\prime}-\chi^{\prime\prime})}\bigg\}\bigg)$
$\times\delta\bigg(\text{sgn}(\tan(\chi^{\prime\prime}))\tan(\chi^{\prime\prime})-\sqrt{\frac{E\cos^{2}(\chi^{\prime}-\chi^{\prime\prime})-H\cos^{2}(\chi^{\prime\prime})}{H\cos^{2}(\chi^{\prime\prime})}}\bigg).$ (B.8)
The Dirac delta constraint in Eq. (B.8) is equivalent to the condition stating that
$\cos^{2}(\chi^{\prime}-\chi^{\prime\prime})=\frac{H}{E}.$ (B.9)
Observe that formula (B.9) treated as an equation with respect to $\chi^{\prime}$ has solutions belonging to the interval $(-\chi,\,-\chi+2\pi)$ only if $H<E$. In order to deduce them, for $H<E$ we build the following set
$\mathcal{A}:=\bigg\{\arccos\sqrt{\frac{H}{E}},\,\pi-\arccos\sqrt{\frac{H}{E}},\,\pi+\arccos\sqrt{\frac{H}{E}},\,2\pi-\arccos\sqrt{\frac{H}{E}}\bigg\}.$ (B.10)
Then, there are four solutions of Eq. (B.9) with respect to $\chi^{\prime}$
$\chi^{\prime}_{0}=\chi^{\prime\prime}+\alpha+2n\pi,$ (B.11)
where $\chi^{\prime\prime}\in(-\chi,\,-\chi+2\pi)$ and $\alpha\in\mathcal{A}$. The value of $n\in\mathbb{Z}$ has been chosen in the way ensuring that $\chi^{\prime}_{0}\in(-\chi,\,-\chi+2\pi)$.
One can show that for the Dirac delta constraint
$F(\chi^{\prime}):=\text{sgn}(\tan(\chi^{\prime\prime}))\tan(\chi^{\prime\prime})-\sqrt{\frac{E\cos^{2}(\chi^{\prime}-\chi^{\prime\prime})-H\cos^{2}(\chi^{\prime\prime})}{H\cos^{2}(\chi^{\prime\prime})}}$
we receive
$\bigg|\frac{dF(\chi^{\prime})}{d\chi^{\prime}}\bigg|^{-1}_{\chi^{\prime}=\chi^{\prime}_{0}}=\frac{\sqrt{H}\sin^{2}(\chi^{\prime\prime})}{\sqrt{\tan^{2}(\chi^{\prime\prime})(E-H)}}=\sqrt{\frac{H}{E-H}}\frac{|\cos(\chi^{\prime\prime})|}{|\sin(\chi^{\prime\prime})|}\sin^{2}(\chi^{\prime\prime}).$ (B.12)
Observe also that for every $\chi^{\prime}_{0}$ the Heaviside function $Y\left(E-H\frac{\cos^{2}(\chi^{\prime\prime})}{\cos^{2}(\chi^{\prime}_{0}-\chi^{\prime\prime})}\right)=1$. Therefore using the identity (B.3) to the solutions (B.11) and applying the determinant (B.12) we obtain that
$\big(W_{Emm^{\prime}}\star_{(\text{M})}W_{\widetilde{E}\widetilde{m}\widetilde{m}^{\prime}}\big)(T,\,\chi,\,H,\,L)=\frac{N_{Emm^{\prime}}N_{\widetilde{E}\widetilde{m}\widetilde{m}^{\prime}}}{2\pi\hbar}e^{i(m-m^{\prime}+\widetilde{m}-\widetilde{m}^{\prime})\chi}$
$\times Y(E-H)\frac{1}{\sqrt{H(E-H)}}\sum_{\alpha\in\mathcal{A}}\int^{-\chi+2\pi}_{-\chi}\text{d}\chi^{\prime\prime}\int^{\infty}_{-\infty}\text{d}L^{\prime\prime}\frac{|\cos(\chi^{\prime\prime})|}{|\sin(\chi^{\prime\prime})|}\sin^{2}(\chi^{\prime\prime})$
$\times\exp\bigg(\frac{2i}{\hbar}\bigg\{-L\tan(\alpha)+L^{\prime\prime}\tan(\chi^{\prime\prime}+\alpha)\bigg\}\bigg)e^{i(m-m^{\prime})(\chi^{\prime\prime}+\alpha)}e^{i(\widetilde{m}-\widetilde{m}^{\prime})\chi^{\prime\prime}}$
$\times Y\left(\widetilde{E}-H\frac{\cos^{2}(\chi^{\prime\prime}+\alpha)}{\cos^{2}(\alpha)}\right)Y\bigg(\frac{\cos(\chi^{\prime\prime}+\alpha)}{\cos(\alpha)}\bigg)Y\bigg(\frac{\cos(\chi^{\prime\prime})}{\cos(\alpha)}\bigg)$
$\times\frac{\cos^{2}(\alpha)}{\cos(\chi^{\prime\prime}+\alpha)\cos(\chi^{\prime\prime})}\frac{1}{\sqrt{\,E\cos^{2}(\alpha)-H\cos^{2}(\chi^{\prime\prime})}\sqrt{\,\widetilde{E}\cos^{2}(\alpha)-H\cos^{2}(\chi^{\prime\prime}+\alpha)}}$
$\times\cos\left(\frac{2L^{\prime\prime}}{\hbar}\sqrt{\frac{\widetilde{E}\cos^{2}(\alpha)-H\cos^{2}(\chi^{\prime\prime}+\alpha)}{H\cos^{2}(\chi^{\prime\prime}+\alpha)}}-(\widetilde{m}+\widetilde{m}^{\prime})\,\arccos\left(\sqrt{\frac{H}{\widetilde{E}}}\frac{\cos(\chi^{\prime\prime}+\alpha)}{\cos(\alpha)}\right)\right)$
$\times\exp\bigg(-i\,\text{sgn}(\tan(\chi^{\prime\prime}))(m+m^{\prime})\,\arccos\bigg\{\sqrt{\frac{H}{E}}\frac{\cos(\chi^{\prime\prime})}{\cos(\alpha)}\bigg\}\bigg).$ (B.13)
Using the property (B.7) again, we integrate over $L^{\prime\prime}$ and obtain
$\big(W_{Emm^{\prime}}\star_{(\text{M})}W_{\widetilde{E}\widetilde{m}\widetilde{m}^{\prime}}\big)(T,\,\chi,\,H,\,L)=\frac{N_{Emm^{\prime}}N_{\widetilde{E}\widetilde{m}\widetilde{m}^{\prime}}}{4}e^{i(m-m^{\prime}+\widetilde{m}-\widetilde{m}^{\prime})\chi}$
$\times Y(E-H)\frac{1}{\sqrt{H(E-H)}}\sum_{\alpha\in\mathcal{A}}\int^{-\chi+2\pi}_{-\chi}\text{d}\chi^{\prime\prime}\frac{|\cos(\chi^{\prime\prime})|}{|\sin(\chi^{\prime\prime})|}\sin^{2}(\chi^{\prime\prime})$
$\times\exp\bigg(-\frac{2i}{\hbar}L\tan(\alpha)\bigg)e^{i(m-m^{\prime})(\chi^{\prime\prime}+\alpha)}e^{i(\widetilde{m}-\widetilde{m}^{\prime})\chi^{\prime\prime}}$
$\times Y\left(\widetilde{E}-H\frac{\cos^{2}(\chi^{\prime\prime}+\alpha)}{\cos^{2}(\alpha)}\right)Y\bigg(\frac{\cos(\chi^{\prime\prime}+\alpha)}{\cos(\alpha)}\bigg)Y\bigg(\frac{\cos(\chi^{\prime\prime})}{\cos(\alpha)}\bigg)$
$\times\frac{\cos^{2}(\alpha)}{\cos(\chi^{\prime\prime}+\alpha)\cos(\chi^{\prime\prime})}\frac{1}{\sqrt{\,E\cos^{2}(\alpha)-H\cos^{2}(\chi^{\prime\prime})}\sqrt{\,\widetilde{E}\cos^{2}(\alpha)-H\cos^{2}(\chi^{\prime\prime}+\alpha)}}$
$\times\exp\bigg(-i\,\text{sgn}(\tan(\chi^{\prime\prime}))(m+m^{\prime})\,\arccos\bigg\{\sqrt{\frac{H}{E}}\frac{\cos(\chi^{\prime\prime})}{\cos(\alpha)}\bigg\}\bigg)$
$\times\exp\bigg(i\,\text{sgn}(\tan(\chi^{\prime\prime}+\alpha))(\widetilde{m}+\widetilde{m}^{\prime})\,\arccos\bigg\{\sqrt{\frac{H}{\widetilde{E}}}\frac{\cos(\chi^{\prime\prime}+\alpha)}{\cos(\alpha)}\bigg\}\bigg)$
$\times\delta\bigg(\text{sgn}(\tan(\chi^{\prime\prime}+\alpha))\tan(\chi^{\prime\prime}+\alpha)-\sqrt{\frac{\widetilde{E}\cos^{2}(\alpha)-H\cos^{2}(\chi^{\prime\prime}+\alpha)}{H\cos^{2}(\chi^{\prime\prime}+\alpha)}}\bigg).$ (B.14)
The Dirac delta constraint in Eq. (B.14) is independent from $\chi^{\prime\prime}$ and leads to the conclusion that $\widetilde{E}=E$. Therefore
$\delta\bigg(\text{sgn}(\tan(\chi^{\prime\prime}+\alpha))\tan(\chi^{\prime\prime}+\alpha)-\sqrt{\frac{\widetilde{E}\cos^{2}(\alpha)-H\cos^{2}(\chi^{\prime\prime}+\alpha)}{H\cos^{2}(\chi^{\prime\prime}+\alpha)}}\bigg)\sim\delta(\widetilde{E}-E).$
For the function
$G(\widetilde{E}):=\text{sgn}(\tan(\chi^{\prime\prime}+\alpha))\tan(\chi^{\prime\prime}+\alpha)-\sqrt{\frac{\widetilde{E}\cos^{2}(\alpha)-H\cos^{2}(\chi^{\prime\prime}+\alpha)}{H\cos^{2}(\chi^{\prime\prime}+\alpha)}}$
we calculate that
$\bigg|\frac{dG(\widetilde{E})}{d\widetilde{E}}\bigg|^{-1}_{\widetilde{E}=E}=2E\,|\sin\big(\chi^{\prime\prime}+\alpha\big)||\cos\big(\chi^{\prime\prime}+\alpha\big)|.$ (B.15)
Observe that $Y\left(\widetilde{E}-H\frac{\cos^{2}(\chi^{\prime\prime}+\alpha)}{\cos^{2}(\alpha)}\right)=1$ for $E=\widetilde{E}$. Thus
$\big(W_{Emm^{\prime}}\star_{(\text{M})}W_{\widetilde{E}\widetilde{m}\widetilde{m}^{\prime}}\big)(T,\,\chi,\,H,\,L)=\frac{N_{Emm^{\prime}}N_{\widetilde{E}\widetilde{m}\widetilde{m}^{\prime}}}{2\sqrt{H(E-H)}}e^{i(m-m^{\prime}+\widetilde{m}-\widetilde{m}^{\prime})\chi}\delta(E-\widetilde{E})Y(E-H)$
$\times\sum_{\alpha\in\mathcal{A}}e^{i(m-m^{\prime})\alpha}\exp\Big[-\frac{2i}{\hbar}L\tan(\alpha)\Big]\int^{-\chi+2\pi}_{-\chi}\text{d}\chi^{\prime\prime}e^{i(m-m^{\prime}+\widetilde{m}-\widetilde{m}^{\prime})\chi^{\prime\prime}}$
$\times\exp\Big(i\,\text{sgn}(\tan(\chi^{\prime\prime}+\alpha))\,(\widetilde{m}+\widetilde{m}^{\prime})\,\arccos\{|\cos(\chi^{\prime\prime}+\alpha)|\}\Big)$
$\times\exp\Big(-i\,\text{sgn}(\tan(\chi^{\prime\prime}))\,(m+m^{\prime})\,\arccos\{|\cos(\chi^{\prime\prime})|\}\Big)Y\bigg(\frac{\cos(\chi^{\prime\prime}+\alpha)}{\cos(\alpha)}\bigg)Y\bigg(\frac{\cos(\chi^{\prime\prime})}{\cos(\alpha)}\bigg).$ (B.16)
Let us notice that for $\theta\neq(k+\frac{1}{2})\pi,\;\;k\in{\mathbb{Z}}$
$\text{sgn}(\tan\theta)\,\arccos|\cos\theta|=\theta-\pi\,\Big\lfloor\frac{\theta+\pi/2}{\pi}\Big\rfloor,$ (B.17)
where $\lfloor\cdot\rfloor$ denotes the floor function.
Changing the variable $\chi^{\prime\prime}\rightarrow\chi^{\prime\prime}-\chi$ we receive
$\big(W_{Emm^{\prime}}\star_{(\text{M})}W_{\widetilde{E}\widetilde{m}\widetilde{m}^{\prime}}\big)(T,\,\chi,\,H,\,L)=\frac{N_{Emm^{\prime}}N_{\widetilde{E}\widetilde{m}\widetilde{m}^{\prime}}}{2\sqrt{H(E-H)}}\delta(E-\widetilde{E})Y(E-H)e^{i(m+m^{\prime}-\widetilde{m}-\widetilde{m}^{\prime})\chi}$
$\times\sum_{\alpha\in\mathcal{A}}e^{i(m-m^{\prime}+\widetilde{m}+\widetilde{m}^{\prime})\alpha}\exp\Big[-\frac{2i}{\hbar}L\tan(\alpha)\Big]\int^{2\pi}_{0}\text{d}\chi^{\prime\prime}e^{2i(\widetilde{m}-m^{\prime})\chi^{\prime\prime}}$
$\times\exp\bigg(-i(\widetilde{m}+\widetilde{m}^{\prime})\big(\pi\Big\lfloor\frac{\chi^{\prime\prime}-\chi+\alpha+\pi/2}{\pi}\Big\rfloor\big)\bigg)$
$\times\exp\bigg(i(m+m^{\prime})\big(\pi\Big\lfloor\frac{\chi^{\prime\prime}-\chi+\pi/2}{\pi}\Big\rfloor\big)\bigg)$
$\times Y\bigg(\frac{\cos(\chi^{\prime\prime}-\chi+\alpha)}{\cos(\alpha)}\bigg)Y\bigg(\frac{\cos(\chi^{\prime\prime}-\chi)}{\cos(\alpha)}\bigg).$ (B.18)
It is necessary to discuss relations between the angles $\chi^{\prime\prime},\;\chi,\;\alpha$ which follow from the Heaviside functions. One can see that
$\alpha\in[0,\pi/2)$ $\Rightarrow(\chi^{\prime\prime}-\chi){\rm mod}\,2\pi\in[0,\pi/2-\alpha)\cup(3\pi/2,2\pi),$
$\alpha\in(\pi/2,\pi]$ $\Rightarrow(\chi^{\prime\prime}-\chi){\rm mod}\,2\pi\in(\pi/2,3\pi/2-\alpha),$
$\alpha\in(\pi,3\pi/2)$ $\Rightarrow(\chi^{\prime\prime}-\chi){\rm mod}\,2\pi\in(5\pi/2-\alpha,2\pi),$
$\alpha\in(3\pi/2,2\pi)$ $\Rightarrow(\chi^{\prime\prime}-\chi){\rm mod}\,2\pi\in[0,\pi/2)\cup(7\pi/2-\alpha,2\pi).$
In the cases first and fourth ones the values of the functions $\Big\lfloor\frac{\chi^{\prime\prime}-\chi+\alpha+\pi/2}{\pi}\Big\rfloor$ and $\Big\lfloor\frac{\chi^{\prime\prime}-\chi+\pi/2}{\pi}\Big\rfloor$ are both even numbers. In the cases second and third ones these values are odd.
Let us assume $m,\,m^{\prime},\,\widetilde{m},\,\widetilde{m}^{\prime}\in\mathbb{Z}$. Thus if $m^{\prime}=\widetilde{m}$, then
$\big(W_{Emm^{\prime}}\star_{(\text{M})}W_{\widetilde{E}m^{\prime}\widetilde{m}^{\prime}}\big)(T,\,\chi,\,H,\,L)=\frac{N_{Emm^{\prime}}N_{\widetilde{E}m^{\prime}\widetilde{m}^{\prime}}}{2\sqrt{H(E-H)}}e^{i(m-\widetilde{m}^{\prime})\chi}\delta(E-\widetilde{E})Y(E-H)$
$\times\sum_{\alpha\in\mathcal{A}}\exp\Big[-\frac{2i}{\hbar}L\tan(\alpha)\Big]e^{i(m+\widetilde{m}^{\prime})\alpha}|\pi-\alpha|(-1)^{(1-\text{sgn}(\cos\alpha))\,(m+\widetilde{m}^{\prime})/2}.$
In order to calculate the sum with respect to $\alpha\in\mathcal{A}$ we use the trigonometric identity $\tan\big(\arccos(x)\big)=\frac{\sqrt{1-x^{2}}}{x}.$ After some manipulations we finally obtain
$\big(W_{Emm^{\prime}}\star_{(\text{M})}W_{\widetilde{E}m^{\prime}\widetilde{m}^{\prime}}\big)(T,\,\chi,\,H,\,L)=\pi N_{Emm^{\prime}}N_{\widetilde{E}m^{\prime}\widetilde{m}^{\prime}}\frac{1}{\sqrt{H(E-H)}}\delta(E-\widetilde{E})Y(E-H)$
$\times e^{i(m-\widetilde{m}^{\prime})\chi}\cos\left(\frac{2L}{\hbar}\sqrt{\frac{E-H}{H}}-(m+\widetilde{m}^{\prime})\,\arccos\sqrt{\frac{H}{E}}\right)$
$=\frac{\pi N_{Emm^{\prime}}N_{\widetilde{E}m^{\prime}\widetilde{m}^{\prime}}}{N_{Em\widetilde{m}^{\prime}}}\delta(E-\widetilde{E})W_{Em\widetilde{m}^{\prime}}(T,\,\chi,\,H,\,L).$ (B.19)
If we denote $W_{Em}:=W_{Emm}$, then for $m=m^{\prime}=\widetilde{m}^{\prime}$ Eq. (B.19) becomes
$(W_{Em}\star_{(\text{M})}W_{\widetilde{E}m}\big)(T,\,\chi,\,H,\,L)=\pi N_{Em}\delta(E-\widetilde{E})W_{Em}(T,\,\chi,\,H,\,L).$ (B.20)
Appendix C Calculation of the Moyal product of Wigner functions $(W_{Emm^{\prime}}\star_{(\text{M})}W_{\widetilde{p}_{x0}\widetilde{p}_{y0}})(T,\,\chi,\,H,\,L)$
The Moyal product of functions $(W_{Emm^{\prime}}\star_{(\text{M})}W_{\widetilde{p}_{x0}\widetilde{p}_{y0}})(T,\,\chi,\,H,\,L)$ is given by the integral
$\big(W_{Emm^{\prime}}\star_{(\text{M})}W_{\widetilde{p}_{x0}\widetilde{p}_{y0}}\big)(T,\,\chi,\,H,\,L)=\frac{N_{Emm^{\prime}}N_{\widetilde{p}_{x0}\widetilde{p}_{y0}}}{(\pi\hbar)^{4}}$
$\times\int^{\infty}_{-\infty}\text{d}T^{\prime}\int^{2\pi}_{0}\text{d}\chi^{\prime}\int^{\infty}_{0}\text{d}H^{\prime}\int^{\infty}_{-\infty}\text{d}L^{\prime}\int^{\infty}_{-\infty}\text{d}T^{\prime\prime}\int^{2\pi}_{0}\text{d}\chi^{\prime\prime}\int^{\infty}_{0}\text{d}H^{\prime\prime}\int^{\infty}_{-\infty}\text{d}L^{\prime\prime}$
$\times Y(E-H^{\prime})\frac{1}{\sqrt{H^{\prime}(E-H^{\prime})}}e^{i(m-m^{\prime})\chi^{\prime}}\cos\bigg(\frac{2L^{\prime}}{\hbar}\sqrt{\frac{E-H^{\prime}}{H^{\prime}}}-(m+m^{\prime})\arccos\sqrt{\frac{H^{\prime}}{E}}\bigg)$
$\times\delta(H^{\prime\prime}-\widetilde{E})\delta(\chi^{\prime\prime}-\widetilde{\chi}_{0})\exp\bigg(\frac{2i}{\hbar}\bigg\{2T\bigg[\sqrt{HH^{\prime}}\cos(\chi-\chi^{\prime})-\sqrt{HH^{\prime\prime}}\cos(\chi-\chi^{\prime\prime})\bigg]$
$+2T^{\prime}\bigg[\sqrt{H^{\prime}H^{\prime\prime}}\cos(\chi^{\prime}-\chi^{\prime\prime})-\sqrt{HH^{\prime}}\cos(\chi-\chi^{\prime})\bigg]$
$+2T^{\prime\prime}\bigg[\sqrt{HH^{\prime\prime}}\cos(\chi-\chi^{\prime\prime})-\sqrt{H^{\prime}H^{\prime\prime}}\cos(\chi^{\prime}-\chi^{\prime\prime})\bigg]$
$+L\bigg[\sqrt{\frac{H^{\prime}}{H}}\sin(\chi-\chi^{\prime})-\sqrt{\frac{H^{\prime\prime}}{H}}\sin(\chi-\chi^{\prime\prime})\bigg]$
$+L^{\prime}\bigg[\sqrt{\frac{H}{H^{\prime}}}\sin(\chi-\chi^{\prime})+\sqrt{\frac{H^{\prime\prime}}{H^{\prime}}}\sin(\chi^{\prime}-\chi^{\prime\prime})\bigg]$
$+L^{\prime\prime}\bigg[\sqrt{\frac{H^{\prime}}{H^{\prime\prime}}}\sin(\chi^{\prime}-\chi^{\prime\prime})-\sqrt{\frac{H}{H^{\prime\prime}}}\sin(\chi-\chi^{\prime\prime})\bigg]\bigg\}\bigg).$ (C.1)
We integrate over $H^{\prime\prime}$ and $\chi^{\prime\prime}$ first.
$\big(W_{Emm^{\prime}}\star_{(\text{M})}W_{\widetilde{p}_{x0}\widetilde{p}_{y0}}\big)(T,\,\chi,\,H,\,L)=\frac{N_{Emm^{\prime}}N_{\widetilde{p}_{x0}\widetilde{p}_{y0}}}{(\pi\hbar)^{4}}$
$\times\int^{\infty}_{-\infty}\text{d}T^{\prime}\int^{2\pi}_{0}\text{d}\chi^{\prime}\int^{\infty}_{0}\text{d}H^{\prime}\int^{\infty}_{-\infty}\text{d}L^{\prime}\int^{\infty}_{-\infty}\text{d}T^{\prime\prime}\int^{\infty}_{-\infty}\text{d}L^{\prime\prime}$
$\times Y(E-H^{\prime})\frac{1}{\sqrt{H^{\prime}(E-H^{\prime})}}e^{i(m-m^{\prime})\chi^{\prime}}\cos\bigg(\frac{2L^{\prime}}{\hbar}\sqrt{\frac{E-H^{\prime}}{H^{\prime}}}-(m+m^{\prime})\arccos\sqrt{\frac{H^{\prime}}{E}}\bigg)$
$\times\exp\bigg(\frac{2i}{\hbar}\bigg\{2T\bigg[\sqrt{HH^{\prime}}\cos(\chi-\chi^{\prime})-\sqrt{H\widetilde{E}}\cos(\chi-\widetilde{\chi}_{0})\bigg]$
$+2T^{\prime}\bigg[\sqrt{H^{\prime}\widetilde{E}}\cos(\chi^{\prime}-\widetilde{\chi}_{0})-\sqrt{HH^{\prime}}\cos(\chi-\chi^{\prime})\bigg]$
$+2T^{\prime\prime}\bigg[\sqrt{H\widetilde{E}}\cos(\chi-\widetilde{\chi}_{0})-\sqrt{H^{\prime}\widetilde{E}}\cos(\chi^{\prime}-\widetilde{\chi}_{0})\bigg]$
$+L\bigg[\sqrt{\frac{H^{\prime}}{H}}\sin(\chi-\chi^{\prime})-\sqrt{\frac{\widetilde{E}}{H}}\sin(\chi-\widetilde{\chi}_{0})\bigg]$
$+L^{\prime}\bigg[\sqrt{\frac{H}{H^{\prime}}}\sin(\chi-\chi^{\prime})+\sqrt{\frac{\widetilde{E}}{H^{\prime}}}\sin(\chi^{\prime}-\widetilde{\chi}_{0})\bigg]$
$+L^{\prime\prime}\bigg[\sqrt{\frac{H^{\prime}}{\widetilde{E}}}\sin(\chi^{\prime}-\widetilde{\chi}_{0})-\sqrt{\frac{H}{\widetilde{E}}}\sin(\chi-\widetilde{\chi}_{0})\bigg]\bigg\}\bigg).$ (C.2)
In the next step we integrate over $^{\prime} , T′′T^{\prime\prime}$ and $L^{\prime\prime}$ using lemmas stating that $\int^{\infty}_{-\infty}\text{d}xe^{ikx}=2\pi\delta(k)$ and $\delta(ax)=\delta(x)/|a|.$ Thus
$\big(W_{Emm^{\prime}}\star_{(\text{M})}W_{\widetilde{p}_{x0}\widetilde{p}_{y0}}\big)(T,\,\chi,\,H,\,L)=\frac{N_{Emm^{\prime}}N_{\widetilde{p}_{x0}\widetilde{p}_{y0}}}{4\pi\hbar}\frac{1}{\sqrt{H^{\prime}}}\int^{2\pi}_{0}\text{d}\chi^{\prime}\int^{\infty}_{0}\text{d}H^{\prime}\int^{\infty}_{-\infty}\text{d}L^{\prime}$
$\times Y(E-H^{\prime})\frac{1}{\sqrt{H^{\prime}(E-H^{\prime})}}e^{i(m-m^{\prime})\chi^{\prime}}\cos\bigg(\frac{2L^{\prime}}{\hbar}\sqrt{\frac{E-H^{\prime}}{H^{\prime}}}-(m+m^{\prime})\arccos\sqrt{\frac{H^{\prime}}{E}}\bigg)$
$\times\delta\bigg[\sqrt{\widetilde{E}}\cos(\chi^{\prime}-\widetilde{\chi}_{0})-\sqrt{H}\cos(\chi-\chi^{\prime})\bigg]$
$\times\delta\bigg[\sqrt{H}\cos(\chi-\widetilde{\chi}_{0})-\sqrt{H^{\prime}}\cos(\chi^{\prime}-\widetilde{\chi}_{0})\bigg]$
$\times\delta\bigg[\sqrt{H^{\prime}}\sin(\chi^{\prime}-\widetilde{\chi}_{0})-\sqrt{H}\sin(\chi-\widetilde{\chi}_{0})\bigg]$
$\times\exp\bigg(\frac{2i}{\hbar}\bigg\{2T\bigg[\sqrt{HH^{\prime}}\cos(\chi-\chi^{\prime})-\sqrt{H\widetilde{E}}\cos(\chi-\widetilde{\chi}_{0})\bigg]$
$+L\bigg[\sqrt{\frac{H^{\prime}}{H}}\sin(\chi-\chi^{\prime})-\sqrt{\frac{\widetilde{E}}{H}}\sin(\chi-\widetilde{\chi}_{0})\bigg]$
$+L^{\prime}\bigg[\sqrt{\frac{H}{H^{\prime}}}\sin(\chi-\chi^{\prime})+\sqrt{\frac{\widetilde{E}}{H^{\prime}}}\sin(\chi^{\prime}-\widetilde{\chi}_{0})\bigg].$
The two last Dirac deltas imply that $H^{\prime}=H$ and $\chi^{\prime}=\chi$. Hence, the first Dirac delta constraint has the solution $\widetilde{E}=\frac{H}{\cos^{2}(\chi-\widetilde{\chi}_{0})}$ that exists only if $\cos(\chi-\widetilde{\chi}_{0})>0$.
Thus our product of the Dirac deltas is equal to
$\frac{4\sqrt{H}}{\cos^{2}(\chi-\widetilde{\chi}_{0})}Y\big(\cos(\chi-\widetilde{\chi}_{0})\big)\delta(\chi^{\prime}-\chi)\delta(H^{\prime}-H)\delta\bigg(\widetilde{E}-\frac{H}{\cos^{2}(\chi-\widetilde{\chi}_{0})}\bigg)$
and the calculated product simplifies to the form
$\big(W_{Emm^{\prime}}\star_{(\text{M})}W_{\widetilde{p}_{x0}\widetilde{p}_{y0}}\big)(T,\,\chi,\,H,\,L)=\frac{N_{Emm^{\prime}}N_{\widetilde{p}_{x0}\widetilde{p}_{y0}}}{\pi\hbar}\frac{Y\big(\cos(\chi-\widetilde{\chi}_{0})\big)}{\cos^{2}(\chi-\widetilde{\chi}_{0})}\delta\bigg(\widetilde{E}-\frac{H}{\cos^{2}(\chi-\widetilde{\chi}_{0})}\bigg)$
$\times Y(E-H)\frac{1}{\sqrt{H(E-H)}}e^{i(m-m^{\prime})\chi}\exp\bigg(-\frac{2i}{\hbar}L\tan(\chi-\widetilde{\chi}_{0})\bigg)$
$\times\int^{\infty}_{-\infty}\text{d}L^{\prime}\cos\bigg(\frac{2L^{\prime}}{\hbar}\sqrt{\frac{E-H}{H}}-(m+m^{\prime})\arccos\sqrt{\frac{H}{E}}\bigg)\exp\bigg(\frac{2i}{\hbar}L^{\prime}\tan(\chi-\widetilde{\chi}_{0})\bigg).$
From the identity (B.7) we obtain that the integral with respect to $L^{\prime}$ equals
$\frac{\pi\hbar}{2}\bigg[\exp\bigg(-i(m+m^{\prime})\arccos\sqrt{\frac{H}{E}}\bigg)\bigg]\delta\bigg(\tan(\chi-\widetilde{\chi}_{0})+\sqrt{\frac{E-H}{H}}\bigg)$
$\frac{\pi\hbar}{2}\bigg[\exp\bigg(i(m+m^{\prime})\arccos\sqrt{\frac{H}{E}}\bigg)\bigg]\delta\bigg(\tan(\chi-\widetilde{\chi}_{0})-\sqrt{\frac{E-H}{H}}\bigg).$
Only one of these terms survives. If $\sin(\chi-\widetilde{\chi}_{0})>0$ then the second component of the sum is relevant. If $\sin(\chi-\widetilde{\chi}_{0})<0,$ the first one contributes in the product. However these both cases lead to the identical conclusion that $E=\frac{H}{\cos^{2}(\chi-\widetilde{\chi}_{0})}=\widetilde{E}$. Moreover, if an angle $\alpha\in(\frac{3\pi}{2},2\pi)$ then $\arccos(\cos(\alpha))=2\pi-\alpha$.
Therefore
$\big(W_{Emm^{\prime}}\star_{(\text{M})}W_{\widetilde{p}_{x0}\widetilde{p}_{y0}}\big)(T,\,\chi,\,H,\,L)=N_{Emm^{\prime}}N_{\widetilde{p}_{x0}\widetilde{p}_{y0}}\frac{Y\big(\cos(\chi-\widetilde{\chi}_{0})\big)}{\cos^{2}(\chi-\widetilde{\chi}_{0})}e^{2im\chi}e^{-i(m+m^{\prime})\widetilde{\chi}_{0}}$
$\times\exp\bigg(-\frac{2i}{\hbar}L\tan(\chi-\widetilde{\chi}_{0})\bigg)\delta\bigg(E-\widetilde{E}\bigg)\delta\bigg(E-\frac{H}{\cos^{2}(\chi-\widetilde{\chi}_{0})}\bigg).$ (C.3)